### v1

# Kai Yao    Marc Juarez

Abstract
Model fingerprint detection techniques have emerged as a promising approach for attributing AI-generated images to their source models, but their robustness under adversarial conditions remains largely unexplored. We present the first systematic security evaluation of these techniques, formalizing threat models that encompass both white- and black-box access and two attack goals: fingerprint removal, which erases identifying traces to evade attribution, and fingerprint forgery, which seeks to cause misattribution to a target model. We implement five attack strategies and evaluate 14 representative fingerprinting methods across RGB, frequency, and learned-feature domains on 12 state-of-the-art image generators. Our experiments reveal a pronounced gap between clean and adversarial performance. Removal attacks are highly effective, often achieving success rates above $80\%$ in white-box settings and over $50\%$ under constrained black-box access. While forgery is more challenging than removal, its success significantly varies across targeted models. We also identify a utility–robustness trade-off: methods with the highest attribution accuracy are often vulnerable to attacks. Although some techniques exhibit robustness in specific settings, none achieves high robustness and accuracy across all evaluated threat models. These findings highlight the need for techniques balancing robustness and accuracy, and identify the most promising approaches for advancing this goal.
I Introduction
The widespread adoption of generative AI models for image generation has created an urgent need for content provenance mechanisms. These models—now integrated into creative tools [adobe2023firefly], content platforms [katsamakas2024gcp], and commercial APIs [openai2023dalle3]—can produce images that are almost visually indistinguishable from authentic photographs. While these technological advances enable new forms of creativity, they also come with risks for misinformation, fraud, and abuse, particularly when AI-generated content is passed as genuine [bird2023typology]. The challenge of reliably identifying and tracing the generated images has thus become critical for maintaining trust and accountability in digital media.
A promising avenue to address this challenge builds on a key property of these models: they tend to leave artifacts—subtle but consistent patterns—in the images they produce [Marra18]. Initially, these patterns were studied as indicators that the images were AI-generated, but researchers showed that the patterns vary across different models and may even serve as fingerprints of the models [Marra19a]. These discoveries have led to the development of model fingerprint detection (MFD), a set of techniques that exploit these fingerprints to address two related problems: deepfake detection—determining whether an image is AI-generated—and model attribution—identifying which specific model generated it.
Existing MFD techniques evaluated under various settings achieve over 90% accuracy for both deepfake detection [Marra19a, Nataraj19, Durall20, Dzanic20, Wang20, Corvi23] and model attribution [Marra18, Marra19b, Yu19, Wang20, Guarnera20, Girish21]. If these accuracies were to hold in realistic scenarios, they would prove useful in bridging the accountability gap in current deployments of generative models. Indeed, being able to identify deepfakes would help mitigate the risks of misinformation and scams, while attributing a harmful output to the specific model that generated it would make it possible to flag issues, understand their causes, and prevent similar problems in the future.
However, high detection accuracy alone is not sufficient to guarantee the practical utility of these techniques. If MFD is to be used as a tool to detect model misuse and be a deterrent for malicious actors, as the academic literature suggests [Marra19a, Marra19b], these techniques must also be reliable in adversarial settings. For example, to prevent the destruction of forensic evidence, it is critical that fingerprints cannot be easily removed.
Yet, security evaluations of MFD are scarce and limited. The few existing studies only evaluate a small subset of MFD techniques, consider narrow threat models and attacker goals, or evaluate only simple image transformations, failing to consider strategic adversaries who are aware of existing MFD techniques and actively attempt to circumvent them [goebel2020adversarial, Wesselkamp22]. This lack of security consideration is in stark contrast to alternative techniques that have been designed for the same purpose. Model watermarking methods, for example, take a more active approach by embedding detectable signals in the generated images, rather than relying on naturally occurring artifacts like MFD [boenisch2021systematic]. While watermarking schemes have been developed with a strong emphasis on security, including systematic evaluations of their robustness to removal and forgery attacks [lin2025crack, Jiang2023EvadingWatermark, Hu2024StableSignature, Zhao2024InvisibleRemovable, an2024waves], the robustness of MFD to similar attacks has received little attention in comparison.
In this paper, we address this gap by providing the first systematic security evaluation of MFD techniques for model attribution. Our evaluation framework includes three plausible threat models that capture varying levels of adversarial knowledge. These threat models lead to five attack strategies that we implement and evaluate on a comprehensive set of MFD techniques compiled from the literature, spanning different feature extraction approaches (pixel-domain, frequency-domain, learned representations), detection algorithms (statistical methods, neural networks), and evaluation settings.
We focus on model attribution rather than deepfake detection because it is a harder problem: model attribution does not merely require detecting the presence of model-induced patterns, but to detect architecture-specific differences across these patterns. Therefore, model attribution is at least as hard as deepfake detection and by evaluating MFD techniques in this setting, we evaluate the worst-case scenario for the fingerprinter, allowing us to draw conclusions for the other scenario: if the techniques prove robust against attacks on model attribution, they should remain robust for the simpler task of deepfake detection.
Our threat models consider two attacker goals: fingerprint removal and fingerprint forgery—the two attacks that have been evaluated in the literature on watermarking for deepfake detection. Within the context of model attribution, removal refers to the suppression of the fingerprint signal such that an image is no longer attributed to its true source, enabling malicious actors to disguise images generated through banned or problematic models [bbc_deepnude_2019]; fingerprint forgery aims to alter the fingerprints such that the images are misattributed to different models, potentially allowing adversaries to falsely implicate legitimate providers.
Our findings show that existing fingerprinting techniques are highly vulnerable to removal attacks, even under constrained attacker capabilities, while success of forgery attacks is more limited. The effectiveness of removal attacks is concerning by itself as it severely limits the utility of these techniques for the applications that have been advocated in the literature [Marra19a, Marra19b]. These results highlight the need for more robust MFD techniques and alternative accountability mechanisms to ensure reliable model attribution in adversarial settings.
To summarize, our main contributions are:
  • We conduct a comprehensive review of the literature on MFD methods across three domains: RGB, frequency, and learned representations, and identify gaps in existing security evaluations.
  • We formalize threat models based on three levels of adversarial knowledge for both removal and forgery attacks, and derive five concrete attack strategies for implementation.
  • We design and implement an evaluation framework that executes these attacks, benchmarking 14 representative MFD methods from the literature for a model attribution task across 12 state-of-the-art generative models spanning GANs, VAEs, and diffusion models on FFHQ dataset.
  • We provide an in-depth analysis and discussion of the attack results, and identify promising directions for strengthening the robustness of MFD methods.
II Background and Related Work
We begin by reviewing existing model fingerprint detection (MFD) techniques for image generation models, along with evaluations of their security properties. This overview highlights key research gaps, particularly regarding the robustness of these techniques for model attribution, gaps that our evaluation framework is designed to address.
II-A Model Fingerprint Detection
Model fingerprint detection (MFD) has gained significant attention as a passive approach to identifying images produced by generative models, such as GANs, VAEs, and diffusion models. Unlike watermarking, which embeds signals during generation, MFD relies on detecting inherent artifacts or “fingerprints” left by the model’s processes. These fingerprints can be detected in statistical irregularities of color and frequency patterns, as well as in feature representations.
Following the classification in Song et al. [Song24], we group MFD methods into those extracting features from the RGB pixel domain, frequency domain, learned features through deep learning, and cross-domain approaches. Across these categories, methods have shown strong performance in controlled settings, often achieving near-perfect accuracy on specific datasets, but their generalization to unseen models or post-processed images varies.
Most of the literature focuses on MFD for deepfake detection, but several studies also evaluate the use of MFD fingerprints for model attribution. While deepfake detection is a binary task, model attribution requires discerning subtle, model-specific patterns among similar architectures, making it a more challenging task. Yet, most evaluations of MFD for model attribution report high detection performance.
RGB Pixel Domain. Methods in this category analyze spatial patterns and color statistics in the image’s pixel values, exploiting deviations from natural image distributions introduced by the generative processes. For instance, early work focused on color saturation irregularities, where the use of normalization layers in generative models tends to suppress over- or under-exposed pixels. One of the earliest MFD approaches applied this strategy for deepfake detection, reporting an AUC of 0.70 for fully GAN-generated faces and 0.61 for partially GAN-synthesized faces, evaluated on a benchmark of roughly 10,000 images from models such as CycleGAN [McCloskey18]. Building on this idea, Nataraj et al. use per-channel co-occurrence matrices to capture inter-channel pixel dependencies, achieving over 99% accuracy on more than 56,000 CycleGAN and StarGAN images [Nataraj19]. Later work extended this to cross-channel co-occurrence, improving deepfake detection of high-quality GAN faces on datasets like CelebA [Nowroozi22]. Although effective for deepfake detection, these methods cannot be directly applied to model attribution without modifications.
Frequency Domain. Generative models often introduce spectral anomalies, such as deficiencies in high-frequency components or unnatural power distributions, which frequency-based methods target through transformations like FFT or DCT. For example, approaches using azimuthally averaged power spectral densities from 2D FFTs have demonstrated near-perfect discrimination (close to 100% accuracy) for deepfake detection between GAN outputs and real images on CelebA and LSUN, detecting patterns in high-frequency components for models like DCGAN and CycleGAN [Durall20]. Other methods extract features from Fourier modes, such as decay exponents and magnitudes, achieving high accuracy (up to 99%) for deepfake detection on ProGAN-generated faces [Dzanic20]. Giudice et al. instead rely on DCT analysis, extracting a 63-dimensional fingerprint vector from the means of AC coefficients of image blocks, achieving an AUC above 0.95 for deepfake detection on datasets like FaceForensics++ [Giudice21]. More recent work combines noise residual autocorrelations with radial and angular spectrum averaging, achieving over 90% model attribution accuracy across GANs, VAEs, and diffusion models on diverse benchmarks including FFHQ and ImageNet subsets [Corvi23].
Learned Features Domain. Deep learning approaches automatically extract implicit fingerprints from high-dimensional representations, and have been widely explored for both deepfake detection and model attribution. For deepfake detection, Marra et al. [Marra18] trained CNNs such as InceptionNet v3 and XceptionNet on social media-like images, achieving around 95% accuracy in distinguishing GAN-generated images from real ones across datasets containing CycleGAN and StarGAN outputs. In the model attribution setting, Wang et al. demonstrated that with careful pre- and post-processing and extensive data augmentation, a ResNet-50 trained on a single generator (ProGAN) can generalize effectively to 11 unseen GAN architectures, reaching 99% detection accuracy [Wang20]. Complementary work has extended CNNs with frequency-domain branches, achieving AUC >> 0.98 on FaceForensics++ and CelebA-DF [Frank20, Qian20]. Yu et al. [Yu19] proposed CNN-based methods that learn model-specific fingerprints, reporting up to 98% attribution accuracy on CelebA across models including ProGAN and StyleGAN, while Girish et al. [Girish21] provided one of the few open-world evaluations, achieving 85–90% recall over 20 GAN architectures—including previously unseen models—using an adaptive clustering algorithm. To improve adaptability, Marra et al. proposed incremental learning strategies that enable detectors to handle new architectures without requiring full retraining, achieving AUC >> 0.9 across up to five GAN models [Marra19b]. Taken together, these results highlight that deep feature-based approaches remain among the most effective techniques for both deepfake detection and model attribution.
Cross-Domain Approaches. To improve performance, Song et al. introduced ManiFPT, a cross-domain fingerprinting framework that unifies multiple domains—RGB, frequency, and learned features [Song24]. The framework defines model fingerprints as deviations of generated images from the true data manifold, formalizing the intuition that the artifacts these models leave in the images are the result of imperfections in modeling the real data distribution. This approach enabled improved model attribution across GANs, VAEs, and diffusion models on benchmarks such as CIFAR-10 and CelebA. A subsequent extension incorporated Riemannian geometry to leverage non-Euclidean manifolds, leading to an average gain of about 6% in model attribution accuracy [Song25].
II-B Existing Robustness Evaluations
While MFD methods perform well in benign settings, their robustness against adversarial evasion remains underexplored—particularly for model attribution, where attackers may attempt to remove or forge subtle, model-specific fingerprints. Existing evaluations largely examine resilience to generic image perturbations or targeted fingerprint removal, typically in black-box settings that do not adapt to the detector’s architecture. Moreover, most studies emphasize deepfake detection rather than attribution. Regarding the study of attacks, prior work has almost exclusively focused on fingerprint removal, overlooking forgery attacks in which adversaries mimic legitimate fingerprints to mislead attribution.
Generic Image Perturbations. Most evaluations test MFD robustness against generic image perturbations such as JPEG compression, blurring, resizing, and rotation, which can unintentionally degrade fingerprints. In the spatial pixel domain, Fourier spectrum–based methods have shown that JPEG compression can reduce detection accuracy by 10–20% on ProGAN datasets, primarily due to the loss of high-frequency components [Dzanic20]. DCT-based approaches demonstrate partial robustness, with accuracy declining by only 5–10% under scaling and rotation on FaceForensics++ (original AUC >> 0.95) [Giudice21]. CNN-based detectors are more resilient under moderate JPEG compression, retaining around 90% accuracy on social media-style datasets with CycleGAN images [Marra18]. Spectrum-domain perturbations, such as high-pass filters, have also been evaluated; for example, detectors like Frank et al.’s [Frank20] experienced up to a 30% accuracy drop on CelebA when subjected to such filters [Wesselkamp22]. Overall, black-box evaluations suggest that although some methods tolerate mild perturbations, others exhibit significant performance degradation even without adaptive adversarial strategies.
Method-Specific Attacks. A limited number of studies have investigated adaptive attacks specifically designed against MFD techniques, aiming to suppress fingerprints while maintaining perceptual quality. For co-occurrence-based detectors such as Nataraj et al.’s [Nataraj19], projected gradient descent (PGD) approximations of the co-occurrence operator can adjust pixel statistics to resemble those of real images, reducing detection accuracy from over 99% to nearly 0% in white-box settings on CycleGAN datasets [Goebel20]. Similarly, frequency-targeted attacks, including peak suppression and mean-spectrum subtraction, have been shown to evade detectors like Frank et al.’s [Frank20], lowering accuracy by 50–80% for deepfake detection on ProGAN and StyleGAN images [Wesselkamp22]. These results demonstrate the vulnerability of MFD methods to adaptive adversaries. However, the extent of their susceptibility to adaptive attacks, particularly against forgery and within attribution settings, remains largely underexplored.
II-C Research Gap
Most MFD studies overlook the security of the techniques despite that the authors propose inherently adversarial applications. For example, uses of fingerprinting in forensics must consider attackers who aim to erase the fingerprints.
We identify four major limitations in the few existing security evaluations:
  1. 1.
    These studies evaluate only simple image perturbations, overlooking strategic adversaries with knowledge of the MFD technique. Since adaptive attacks are more likely to succeed, these evaluations give a false sense of security.
  2. 2.
    The only two evaluations that include adaptive attacks evaluate the attacks against specific MFD techniques, instead of using appropriate baselines or evaluating against a comprehensive range of techniques. An evaluation on a single technique does not allow to extrapolate to the effectiveness of the attacks against other techniques.
  3. 3.
    Despite the fact that these studies propose MFD techniques for model attribution, most evaluations only assess security for deepfake detection, leaving security regarding model attribution unaddressed.
  4. 4.
    All prior research has focused on removal attacks, leaving robustness to forgery attacks underexplored. As we explain in Section III-B, although these two attacks are related, effectiveness of one does not imply the other.
Our work closes this gap by formalizing threat models for MFD techniques and systematically deriving attack strategies from these threat models. To provide comprehensiveness, we group MFD techniques with similar characteristics and evaluate them under a common evaluation framework that implements all our variations of removal and forgery attacks.
III Threat Model
To systematically evaluate the security of model fingerprint detection (MFD) techniques for model attribution tasks, we formalize adversary models with varying levels of knowledge and access—ranging from a full black-box setting, where the adversary has no information about the MFD system, to a white-box setting, where the adversary has complete access. Under these different visibility levels, we study two objectives: fingerprint removal and fingerprint forgery. Each access level motivates distinct attack strategies, which we implement and evaluate in later sections.
III-A System Model
Let $G:\mathcal{Z}\rightarrow\mathcal{X}$ denote a generative model that maps latent variables $z\in\mathcal{Z}$ to images $x=G(z)\in\mathcal{X}$. We define a model fingerprint detection (MFD) system for model attribution as
$\mathcal{F}(x)=h(\phi(x)),$ (1)
where $\phi:\mathcal{X}\rightarrow\mathcal{A}$ is the fingerprint extractor that derives fingerprint features from the input image, and $h:\mathcal{A}\rightarrow\mathcal{Y}$ is the model attribution classifier that predicts the source identity $y\in\mathcal{Y}=\{G_{1},G_{2},\ldots,G_{K}\}$ among $K$ candidate generators.
This formulation—decoupling the fingerprint extractor $\phi$ from the attribution classifier hh—matches the structure of most MFD methods in the literature. Some approaches instead employ end-to-end neural networks, where fingerprint extraction is not a separate function (e.g., [Qian20, Wang20]). In such cases, the fingerprint can be defined implicitly: the preceding layers act as the fingerprint extractor $\phi$, while the final classification layer serves as the model attribution classifier $h$.
III-B Adversarial Objectives
The adversary aims to generate an adversarial image $x^{\prime}$ that misleads the MFD system $\mathcal{F}$. We impose a general distance constraint between the original and adversarial images:
$d(x,x^{\prime})\leq\varepsilon,$ (2)
where $d:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{\geq 0}$ is a distance function and $\varepsilon$ specifies the perturbation budget. Typical choices of $d$ include $\ell_{\infty}$ or $\ell_{2}$ norms, LPIPS [zhang2018unreasonable], or PSNR. This constraint ensures that adversarial manipulations do not degrade the quality of the attacks images excessively.
We consider two adversarial objectives under this constraint:
  • Fingerprint Removal. Given an image $x=G(z)$ with true source label $y=G$, the attacker creates $x^{\prime}$ such that
    $\mathcal{F}(x^{\prime})\neq y,\quad d(x,x^{\prime})\leq\varepsilon.$ (3)
    A successful removal attack suppresses enough model-specific information in the fingerprint features to cause model attribution classifier $h$ to misclassify $x^{\prime}$. Since the MFD system operates in a closed world of candidate generators, such misclassification typically results in a low-confidence attribution to a different model.
  • Fingerprint Forgery. Given $x=G(z)$ with true source label $y=G$ and a chosen target label $y_{t}=G_{t}\neq y$, the adversary produces $x^{\prime}$ such that
    $\mathcal{F}(x^{\prime})=y_{t},\quad d(x,x^{\prime})\leq\varepsilon.$ (4)
    Unlike removal, forgery attacks are targeted: the adversary induces a specific misclassification, forcing the MFD system to attribute $x^{\prime}$ to the target generator $G_{t}$ rather than the true source $G$. This requires introducing sufficient features resembling $G_{t}$’s fingerprints so that the model attribution classifier $h$ outputs the target label $y_{t}$.
III-C Adversary Knowledge Levels
We distinguish three levels of adversarial knowledge about the MFD system $\mathcal{F}$, depending on the adversary’s access to the fingerprint extractor $\phi$, the model attribution classifier $h$, and their knowledge of the model attribution task:
  1. 1.
    White-box Access. The adversary has complete access to both $\phi$ and $h$, enabling end-to-end gradient computations through $\mathcal{F}$. This represents the strongest adversary and, as our results show, allows for highly effective removal and forgery attacks. While less common, this setting is realistic when the MFD system is open-sourced or if a malicious insider gains direct access to the target model.
  2. 2.
    Black-box Access I. The adversary has no access to $\phi$ or $h$, but knows the model attribution task, i.e., the candidate generators $\mathcal{Y}$. By sampling images from all $G\in\mathcal{Y}$, the adversary can train a surrogate classifier $h_{s}$ directly on generated images, bypassing fingerprint extraction. Adversarial examples $x^{\prime}$ are then crafted against $h_{s}$ with the expectation that they transfer to the true MFD system $\mathcal{F}$, leveraging adversarial transferability [goodfellow2014explaining]. This scenario is realistic since attribution systems are typically concealed to prevent manipulation. Attackers can only use publicly available generators to build surrogates, not to replicate the defender’s pipeline but to create adversarial examples that may transfer to the MFD system.
  3. 3.
    Black-box Access II. The adversary has no knowledge of $\phi$, $h$, or the model attribution task. Typical low-cost attacks in this setting are generic image manipulations such as compression, noise, or blurring. This setting makes the least assumptions and is also realistic: defenders usually hide system details, and opportunistic attackers may only apply low-cost transformations in the hope of reducing model attribution confidence.
Across all three knowledge levels, we assume the adversary does not have internal access to the target generator $G$ (nor to $G_{t}$ in forgery), but can query any of the candidate generators and sample images from them. This reflects real-world deployments of generative models that expose subscription-based access through user APIs. Internal access to the generator, however, represents a stronger threat model outside our scope: although an attacker compromising model provider infrastructure may obtain internal information about the generator, such access enables threats far more severe than fingerprint forgery or removal attacks, making these attacks moot. In addition, although access to a generator might give the adversary an advantage against an MFD method designed for that specific generator, most MFD methods—like our attacks—are generator-agnostic and applicable across generator families.
Note that removal attacks are feasible under all three knowledge levels, but forgery attacks are infeasible in the Black-box II setting because the adversary lacks access to $G_{t}$ and thus lacks the information needed to cause a targeted misattribution.
IV Attack Implementation
Having formalized the threat models, we now describe the concrete implementations of attack strategies against an MFD system for model attribution. We propose five attack strategies, each motivated by the adversary’s knowledge level and the mathematical properties of the fingerprint extractor $\phi$. These are general strategies, not tied to any single MFD implementation; they apply whenever the stated assumptions about $\phi$ and the adversary’s access hold. This makes our evaluation framework broadly applicable to unseen MFD methods.
IV-A White-box Attack Strategies
Under White-box Access, the adversary has full access to both $\phi$ and $h$. Because $h$ is typically a differentiable neural-network classifier, and $\mathcal{F}$ becomes differentiable whenever $\phi$ is differentiable (or can be made so), the adversary can apply gradient-based optimization to craft adversarial examples [goodfellow2014explaining, madry2017towards]. Concretely, the attacker computes gradients of an attribution loss of $\mathcal{F}$ w.r.t. the input image and iteratively perturbs pixels to reduce the confidence of the correct class (removal) or to increase the confidence of a chosen target class (forgery).
Depending on properties of $\phi$, we distinguish three white-box optimization strategies:
  • W1: Direct Differentiation. When $\phi$ is differentiable, attackers directly apply gradient-based attacks on $h(\phi(x))$.
  • W2: Analytic Approximation ($\tilde{\phi}$). If $\phi$ is non-differentiable but admits a differentiable analytic approximation, the adversary constructs a differentiable surrogate $\tilde{\phi}(x)$ (e.g., soft histograms [Yusuf2020DifferentiableHistogram], relaxed co-occurrence statistics [Levy2014NeuralWordEmbedding]) and optimizes $h(\tilde{\phi}(x))$.
  • W3: Surrogate Extractor ($\phi_{s}$). If $\phi$ is neither differentiable nor amenable to analytic approximation, the adversary trains a differentiable surrogate extractor $\phi_{s}(x)$ using pairs $\{(x,\phi(x))\}$ sampled from generators $G\in\mathcal{Y}$ to approximate $\phi$. Attacks then optimize $h(\phi_{s}(x))$.
Figure 1 depicts which white-box attacks (W1–W3) apply given $\phi$’s properties. These strategies differ in their assumptions: W1 makes the strictest, requiring perfect gradient information. W2 and W3 are possible when there is gradient information but offer no advantage over W1. Similarly, W3 is possible when W2 is feasible, but not the other way around.
Figure 1: Decision tree for selecting a white-box attack strategy. The applicability of white-box attacks depends on whether the fingerprint extractor, $\phi$, is differentiable and, if not, whether it can be analytically approximated.
IV-B Black-box Attack Strategies
When the attacker has limited knowledge of $\phi$ and $h$, attacks rely on surrogate training or generic image manipulations:
  • B1: Surrogate Attribution Classifier ($h_{s}$). Under Black-box Access I, the candidate set $\mathcal{Y}$ is known but $\phi$ is not. The adversary can sample images from each $G\in\mathcal{Y}$ and train a surrogate classifier $h_{s}(x)$ that approximates the mapping $x\mapsto h(\phi(x))$ on the image space. The adversary then applies gradient-based attacks on $h_{s}$, hoping that they transfer to the targeted MFD system. Note that the surrogate $\phi_{s}$ in W3 is a surrogate of the fingerprint extractor $\phi$, while this surrogate, $h_{s}$, is for the entire model attribution classifier $h$.
  • B2: Image Transformations. Under Black-box Access II, without knowledge of $\phi$, $h$, or $\mathcal{Y}$, a typical adversary applies generic image transformations such as JPEG compression, noising, blurring, or resizing. These untargeted manipulations aim to degrade fragile fingerprint traces and reduce attribution confidence and without access to the target generator they cannot achieve forgery.
Under Black-box Access II, adversaries could pursue more sophisticated attacks, such as training a surrogate on outputs from the target generator and a dataset for the negative class. However, these require substantial sampling and optimization, and would still underperform B1 under Black-box Access I (with knowledge of $\mathcal{Y}$). for this reason, we restrict Black-box Access II to low-cost image perturbations as a representative, practical attack strategy.
IV-C Unified Optimization Objective
All gradient-based attacks (W1–W3, B1) can be expressed within a single constrained optimization framework:
$\min_{x^{\prime}}\;\;\mathcal{L}_{\text{adv}}(\mathcal{F}^{\prime}(x^{\prime}))\quad\text{subject to}\quad d(x,x^{\prime})\leq\varepsilon,$ (5)
where the constraint is the same as in Eq. 2. Here, $\mathcal{F}^{\prime}$ denotes the attribution model available to the attacker in each setting:
  • W1: $\mathcal{F}^{\prime}(x)=\mathcal{F}=h(\phi(x))$, i.e., the true MFD system.
  • W2: $\mathcal{F}^{\prime}(x)=h(\tilde{\phi}(x))$, where $\tilde{\phi}$ is an analytic differentiable approximation of $\phi$.
  • W3: $\mathcal{F}^{\prime}(x)=h(\phi_{s}(x))$, where $\phi_{s}$ is a learned surrogate extractor approximating $\phi$.
  • B1: $\mathcal{F}^{\prime}(x)=h_{s}(x)$, where $h_{s}$ is a surrogate attribution classifier trained on images.
The adversarial loss $\mathcal{L}_{\text{adv}}$ is defined according attack objective:
  • Fingerprint Removal: Suppress attribution to the true source $y$ by optimizing a loss that discourages high confidence on that label. We use a cross-entropy–based objective:
    $\mathcal{L}_{\text{adv}}=-\log\left(1-P_{\mathcal{F}^{\prime}}(y\mid x^{\prime})\right),$
    which provides stable gradients and effectively drives the prediction away from $y$.
  • Fingerprint Forgery: Induce attribution to a chosen target generator with label $y_{t}$. We minimize the cross-entropy loss with respect to the target:
    $\mathcal{L}_{\text{adv}}=-\log P_{\mathcal{F}^{\prime}}(y_{t}\mid x^{\prime}),$
    encouraging confident misattribution toward $y_{t}$ and yielding more stable and effective optimization.
Summary. The attack implementation taxonomy detailed in this section provides an extendable framework for evaluating and comparing removal and forgery robustness across MFD techniques and adversary knowledge levels. We will introduce our experimental setup in Section V and empirically evaluate all attack strategies in Section VI.
TABLE I: MFD techniques for generative model attribution. The table shows domain, fingerprint description, size, original attribution accuracy, mathematical properties (differentiability $\nabla_{x}\mathcal{\phi}$ and analytic approximation $\tilde{\phi}$), and attack compatibility (W1: direct gradient attacks, W2: analytic approximation attacks, W3: surrogate extractor attacks, B1-B2: black-box attacks).
Method Domain Fingerprint Description Size Accuracy (% $\pm$ Std) $\nabla_{x}\mathcal{\phi}$ $\tilde{\phi}$ W1 W2 W3 B1 B2
McCloskey18 [McCloskey18] RGB Saturation pixel proportions 1×4 23.81 $\pm$ 0.94 - $\bullet$ $\bullet$ $\bullet$ $\bullet$
Nataraj19 [Nataraj19] RGB Co-occurrence matrices 3×256×256 90.96 $\pm$ 0.74 - $\bullet$ $\bullet$ $\bullet$ $\bullet$
Nowroozi22 [Nowroozi22] RGB Cross-band co-occurrence 6×256×256 87.31 $\pm$ 1.15 - $\bullet$ $\bullet$ $\bullet$ $\bullet$
Song24-RGB [Song24] RGB RGB manifold deviations 3×256×256 63.82 $\pm$ 0.76 - - $\bullet$ $\bullet$ $\bullet$
Marra19a [Marra19a] Frequency Denoising residuals 3×256×256 66.67 $\pm$ 2.16 - $\bullet$ $\bullet$ $\bullet$ $\bullet$
Durall20 [Durall20] Frequency Radial power spectrum 1×179 80.90 $\pm$ 1.02 - $\bullet$ $\bullet$ $\bullet$ $\bullet$
Dzanic20 [Dzanic20] Frequency Power law decay params 1×3 78.98 $\pm$ 0.11 - $\bullet$ $\bullet$ $\bullet$ $\bullet$
Giudice21 [Giudice21] Frequency DCT coefficient statistics 1×63 94.88 $\pm$ 1.04 - $\bullet$ $\bullet$ $\bullet$
Corvi23-R [Corvi23] Frequency Noise residual autocorr 1×4225 39.65 $\pm$ 1.19 - $\bullet$ $\bullet$ $\bullet$
Corvi23-S [Corvi23] Frequency Radial/angular spectra 1×144 27.11 $\pm$ 8.62 - $\bullet$ $\bullet$ $\bullet$
Song24-Freq [Song24] Frequency Frequency manifold deviations 1×256×256 84.21 $\pm$ 0.85 - - $\bullet$ $\bullet$ $\bullet$
Qian20 [Qian20] Learned F3-Net frequency features - 90.12 $\pm$ 0.84 - $\bullet$ $\bullet$ $\bullet$
Wang20 [Wang20] Learned ResNet50 features - 98.47 $\pm$ 0.51 - $\bullet$ $\bullet$ $\bullet$
Song24-SL [Song24] Learned SL manifold deviations 1×2048 87.90 $\pm$ 0.91 - - $\bullet$ $\bullet$ $\bullet$
V Experimental Setup
We now describe the experimental setup used to evaluate the attack strategies, which includes the generative models, MFD methods, and datasets selected for evaluation. We also describe the training of the attribution models and the details of our implementations of the attack strategies.
V-A Image Generators
We evaluate our framework on a model attribution task involving images from 12 widely used state-of-the-art generative models, covering 3 major architectural classes: 6 GAN models (StyleGAN2 [Karras2020ADA], StyleGAN3 [Karras2021StyleGAN3], GANformer [Hudson2021GANformer], StyleSwin [Zhang2022StyleSwin], R3GAN [Huang2024R3GAN], and CIPS [Anokhin2021CIPS]), 3 VAE models (VDVAE [Child2021VDVAE], VQVAE [van2017neural], and NVAE [vahdat2020nvae]), and 3 diffusion models (NCSN++ [Song2021ScoreSDE], LDM [rombach2022high], and ADM [dhariwal2021diffusion]). All models are pre-trained on the FFHQ dataset and generate 256×256 RGB images. All models are sourced from official repositories with pre-trained checkpoints to ensure reproducibility.
V-B MFD Methods
We evaluate 14 MFD methods spanning the RGB, frequency, and learned feature domains. Although this set is not exhaustive, it is representative: it includes influential methods that have been widely studied for model attribution (e.g., those covered in Song et al. [Song24]) and thus reflects the performance and robustness trends of the state of the art.
Table I summarizes their key characteristics, including the feature domain (RGB, frequency, or learned representations [Song24]), a brief description of the fingerprint features, and the fingerprint size (i.e., the output of $\phi$, or - for end-to-end methods). We also report attribution accuracy of the MFD methods, obtained by reproducing their evaluations on the new pool of 12 candidate generative models.
The remaining columns indicate whether the method is differentiable, whether it admits an analytic approximation, and which attack strategies are applicable. For non-differentiable methods that allow an analytic approximation, we introduce differentiable variants $\tilde{\phi}$ to enable W2 attacks.
V-C Models and Training Details
We train three neural network models to evaluate fingerprint robustness across different attack settings. All models are optimized with Adam for 100 epochs on $1,000$ images/fingerprints per generative model, using a batch size of 32, a validation split of 20%, and early stopping to prevent overfitting. The learning rate is fixed at $1\times 10^{-4}$.
The models that require training are:
  • Attribution Classifier ($h$). This model maps fingerprint features to source model labels and serves as the primary target for attacks. It is implemented as a MLP with hidden dimensions [512, 256, 128] and ReLU activations, trained with cross-entropy loss. The architecture is fixed across all MFD methods, while the input dimensionality depends on the fingerprint size of the specific method.
  • Surrogate Attribution Classifier ($h_{s}$). This model approximates $h$ when fingerprint features are unavailable, enabling black-box attacks (B1). It takes images as input and predicts source model labels. The architecture is a CNN with four convolutional blocks (64→128→256→512 channels), followed by fully connected layers, trained with cross-entropy loss.
  • Surrogate Fingerprint Extractor ($\phi_{s}$). This model approximates non-differentiable fingerprint extractors by regressing from images to fingerprint features. It uses a CNN with four convolutional blocks (64→128→256→512 channels) and fully connected layers, trained with MSE loss to match the dimensionality of the original extractor’s outputs.
V-D Attack Implementation and Evaluation
We implement five adversarial attack strategies across the white-box (W1–W3) and black-box (B1–B2) threat models detailed in Section IV to evaluate fingerprint robustness. While removal attack are not targeted, forgery attacks require the attacker to induce a misclassification towards a specific target candidate generator label different from the true label. In the evaluation, we choose the target label uniformly at random on each execution of a forgery attack.
For gradient-based attacks (W1–W3, B1), we employ Projected Gradient Descent (PGD) [madry2017towards] with momentum and adaptive step size selection. We randomly sample 100 images from each generative model, and conduct attacks only on the images that are accurately classified by the true attribution model $h$. The rationale is that the attacker gains no advantage from attacking images that are already misclassified.
For each attacked image, we evaluate PGD with step sizes in $\{0.0005,0.001,0.005,0.01,0.05,0.1\}$ and consider the attack successful if any of the step sizes succeeds. The attacker can also do that in practice because they have access to images and MFD methods, so the attacker can tune the hyperparameters of their attacks for each fingerprinting technique. We then aggregate successful attempts across all images to compute the Attack Success Rate (ASR), which estimates the probability that the attack strategy will successfully remove or forge fingerprints in any given image.
PGD is limited to 50 iterations, optimizing the cross-entropy loss with a momentum coefficient of $0.9$, random initialization, and gradient clipping. The perturbation is bounded by $\epsilon=0.025$ (i.e., $\ell_{\infty}$ norm). We adopt the $\ell_{\infty}$ norm for its computational efficiency and its direct interpretation as a pixel value bound, while also verifying that perturbations remain visually acceptable. Across all attacks, this $\epsilon$ level consistently yields LPIPS <0.05<0.05 and PSNR >35>35 dB, which indicates satisfactory perceptual quality [zhang2018lpips, huynh2008psnr] (see Appendix A for more details). We also report the relationship between $\epsilon$ and ASR and LPIPS/PSNR in Figure 8 in Appendix D.
All gradient-based attacks support both removal and forgery objectives. For non-gradient attacks (B2), we employ four standard image perturbations with parameter settings chosen to preserve high image quality comparable to gradient-based attacks (i.e. LPIPS <0.05<0.05 and PSNR >35>35 dB). Specifically, we apply additive Gaussian noise with $\sigma=0.005$ (B2-Noise), Gaussian blur with $\sigma=0.5$ using a $3{\times}3$ kernel (B2-Blur), JPEG compression at quality $95$ (B2-JPEG), and isotropic downsampling to $0.9\times$ followed by upsampling (B2-Resize). Each perturbation is applied once independently to the original image, simulating one single adversarial query. We also conduct an ablation over these hyperparameters and report their impact on attack success in Figure 9 in Appendix D. Since B2’s perturbations are untargeted, we evaluate them only for removal attacks.
We conduct all experiments with PyTorch on NVIDIA A100 GPUs. The reported results are the mean and standard deviation over $5$ independent runs, where each run involves training all models in the pipeline and then conducting an attack. Upon acceptance, we will publish all our source code with documentation on how to reproduce the experiments.
VI Results
This section presents the results of our security evaluation. We organize it into five parts: first, we assess MFD technique robustness to removal across threat models and attack strategies; second, we repeat the evaluation for forgery; third, we study the relationship between removal and forgery; fourth, we test whether higher utility of the MFD technique is an indicator of vulnerability to the attacks; and, fifth, we investigate the impact of model’s type and architecture on ASR.
Overall, our evaluation reveals that existing MFD techniques are highly susceptible to adversarial manipulation. Removal attacks are particularly successful; critically, even simple B2 perturbations achieve ASR $50\%$ for many MFD methods. The factor that dominates attack success is the MFD technique rather than the fingerprinted model, with more accurate techniques being slightly more vulnerable to the attacks.
VI-A Robustness to Removal
TABLE II: Attack success rates (ASR, mean $\pm$ std, in %) of removal attacks against evaluated MFD methods. White-box attacks include direct gradient-based optimization (W1), analytic approximations (W2), and surrogate fingerprint extractors (W3). Black-box attacks include surrogate classifiers (B1) and fingerprint-agnostic perturbations (B2: Gaussian noise, blur, JPEG compression, resizing). Higher ASR indicates greater vulnerability. The maximum values per method are highlighted in bold. The row separators indicate the domains in Table I: RGB, Frequency, and Learned.
Method W1 W2 W3 B1 B2-Noise B2-Blur B2-JPEG B2-Resize
McCloskey18 [McCloskey18] 36.70 $\pm$ 5.16 59.36 $\pm$ 28.81 13.49 $\pm$ 0.55 2.09 $\pm$ 1.00 13.33 $\pm$ 5.70 2.88 $\pm$ 1.75 11.97 $\pm$ 5.89
Nataraj19 [Nataraj19] 97.53 $\pm$ 1.84 86.54 $\pm$ 12.49 84.75 $\pm$ 0.94 27.97 $\pm$ 0.85 38.31 $\pm$ 3.16 46.16 $\pm$ 4.09 51.42 $\pm$ 1.14
Nowroozi22 [Nowroozi22] 84.94 $\pm$ 1.32 67.87 $\pm$ 1.75 84.55 $\pm$ 0.97 22.47 $\pm$ 1.69 29.00 $\pm$ 1.26 32.43 $\pm$ 1.66 42.39 $\pm$ 2.87
Song24-RGB [Song24] 91.38 $\pm$ 2.63 2.34 $\pm$ 0.61 0.00 $\pm$ 0.00 1.19 $\pm$ 0.40 0.18 $\pm$ 0.32 0.00 $\pm$ 0.00
Marra19a [Marra19a] 100.00 $\pm$ 0.00 40.71 $\pm$ 5.20 9.55 $\pm$ 2.10 1.67 $\pm$ 1.07 13.50 $\pm$ 9.26 4.10 $\pm$ 1.06 5.04 $\pm$ 1.18
Durall20 [Durall20] 99.90 $\pm$ 0.20 95.23 $\pm$ 7.93 99.46 $\pm$ 0.94 20.83 $\pm$ 5.80 61.63 $\pm$ 2.21 39.04 $\pm$ 4.51 94.12 $\pm$ 1.11
Dzanic20 [Dzanic20] 98.90 $\pm$ 1.71 99.10 $\pm$ 1.37 99.78 $\pm$ 0.22 40.33 $\pm$ 0.18 90.59 $\pm$ 4.08 48.26 $\pm$ 0.54 98.63 $\pm$ 0.61
Giudice21 [Giudice21] 100.00 $\pm$ 0.00 92.74 $\pm$ 8.01 21.99 $\pm$ 1.30 59.21 $\pm$ 3.61 45.48 $\pm$ 1.80 81.18 $\pm$ 1.76
Corvi23-R [Corvi23] 100.00 $\pm$ 0.00 35.16 $\pm$ 3.84 2.44 $\pm$ 1.03 17.02 $\pm$ 4.26 1.11 $\pm$ 0.37 25.20 $\pm$ 6.53
Corvi23-S [Corvi23] 77.44 $\pm$ 27.47 15.08 $\pm$ 7.42 0.42 $\pm$ 0.72 27.90 $\pm$ 18.07 0.42 $\pm$ 0.72 32.21 $\pm$ 15.68
Song24-Freq [Song24] 31.08 $\pm$ 3.82 70.56 $\pm$ 6.94 5.30 $\pm$ 0.87 50.55 $\pm$ 3.78 6.79 $\pm$ 1.55 68.64 $\pm$ 0.63
Qian20 [Qian20] 90.71 $\pm$ 6.43 82.05 $\pm$ 6.12 56.08 $\pm$ 38.06 67.12 $\pm$ 28.49 26.28 $\pm$ 6.65 75.90 $\pm$ 20.89
Wang20 [Wang20] 100.00 $\pm$ 0.00 78.19 $\pm$ 18.90 0.67 $\pm$ 0.77 37.08 $\pm$ 54.53 34.17 $\pm$ 57.01 49.04 $\pm$ 44.23
Song24-SL [Song24] 84.45 $\pm$ 0.85 55.86 $\pm$ 1.97 2.68 $\pm$ 0.74 20.18 $\pm$ 1.66 7.77 $\pm$ 1.51 34.61 $\pm$ 2.59
We evaluate all MFD methods against their compatible attack strategies, as summarized in Table I. ASRs are reported separately for removal (Table II) and forgery (Table III) attacks, with the highest ASR per method shown in bold.
Removal attacks are broadly effective across most MFD methods. In many cases, the strongest attack achieves over $80\%$ ASR under the $\ell_{\infty}$ norm constraint, indicating that fingerprints can often be suppressed while maintaining high perceptual quality. Corresponding LPIPS and PSNR values remain within good quality ranges for most attacks, and representative attacked images confirm that the perturbations are imperceptible (see Tables IV, V, and Figure 5 in Appendix A).
The main exceptions are Marra19a and Song24-RGB, which exhibit strong resistance to most attacks unless white-box access is available. McCloskey18, Corvi23-R, and Corvi23-S also show some resistance to black-box attacks; however, their baseline (non-adversarial) accuracies are substantially lower. Marra19a relies on noise residuals (image minus its BM3D-denoised version), and our results indicate that structured residuals are difficult to fully eliminate under black-box settings, making the method more robust. Song24-RGB measures the distance of a sample to the true data manifold in the RGB domain. Because fingerprints from the same model likely form tight clusters in this space and our attacks enforce an $\ell_{\infty}$ perturbation bound in the RGB domain, perturbed samples rarely escape their cluster, limiting the effectiveness of removal. In contrast, Song24-Freq and Song24-SL lack such constraints in the frequency or learned feature domains, which explains their greater vulnerability to black-box attacks compared to Song24-RGB. Since these methods are highly vulnerable to white-box attacks yet robust in black-box scenarios, it is likely that their fingerprints are inherently less transferable from the image space. Taken together, these findings suggest that noise residual- and manifold-based representations holds promise for robustness against removal in black-box scenarios.
Table II shows that ASR varies based on adversarial knowledge. White-box attacks (W1–W3) generally outperform black-box attacks (B1 and B2) across MFD techniques, which is expected since white-box attackers have privileged visibility into the MFD system. Among white-box, W1 attacks consistently achieve the best performance—when they are applicable. However, W1 attacks are not applicable for techniques that do not support direct gradient optimization. For these cases, W2 and W3 attacks rely on approximations of the gradients. As seen in the table, these approximations sometimes underperform relative to B1 strategies, which are also approximated. This counterintuitive result arises when the approximation errors of W2 and W3 exceed those of B1.
Even when B1 is not the optimal strategy, its success remains competitive with white-box approaches for several MFD methods, showing that black-box surrogate-based attacks can achieve relatively high success rates with only knowledge of the MFD technique implementation. In many applications, such as forensic use cases, MFD techniques cannot realistically be kept secret, making B1 removal attacks a practical threat.
Although adaptive strategies (W1–W3, and B1) always outperform fingerprint-agnostic strategies based on simple image manipulations (B2), these generic black-box perturbations alone often achieve ASRs $50\%$ and achieve ASR $90\%$ for one of the techniques (Dzanic20), indicating that fingerprint traces can be significantly disrupted without any knowledge of the underlying MFD technique. This further demonstrates the fragility of current MFD methods and heightens concerns about the applicability of these techniques in adversarial settings where not even costly attacks are needed.
ASRs show low standard deviation across most attack instances. However, some attacks exhibit high variance, with W3 against McCloskey18 showing up to $28.81$ pp. difference in ASR. The MFD technique appears to be the primary driver of this variance, creating uncertainty for users about attack robustness. This variability suggests certain samples are inherently more vulnerable to fingerprint removal. We leave a detailed investigation of the underlying causes for future work.
TABLE III: Attack success rates (ASR, mean $\pm$ std, in %) of forgery attacks. White-box attacks include direct gradient-based optimization (W1), analytic approximations (W2), and surrogate fingerprint extractors (W3). Black-box attacks include surrogate classifiers (B1). The maximum value per method is highlighted in bold.
Method W1 W2 W3 B1
McCloskey18 [McCloskey18] 7.51 $\pm$ 2.82 10.43 $\pm$ 2.81 2.10 $\pm$ 0.06
Nataraj19 [Nataraj19] 51.64 $\pm$ 2.78 9.85 $\pm$ 3.17 11.27 $\pm$ 0.72
Nowroozi22 [Nowroozi22] 12.09 $\pm$ 3.57 6.84 $\pm$ 0.73 10.20 $\pm$ 1.57
Song24-RGB [Song24] 16.42 $\pm$ 4.59 0.00 $\pm$ 0.00
Marra19a [Marra19a] 97.85 $\pm$ 2.90 4.46 $\pm$ 1.28 0.50 $\pm$ 0.52
Durall20 [Durall20] 57.67 $\pm$ 8.63 23.49 $\pm$ 2.70 18.18 $\pm$ 1.13
Dzanic20 [Dzanic20] 15.19 $\pm$ 5.09 25.53 $\pm$ 9.61 25.51 $\pm$ 3.93
Giudice21 [Giudice21] 98.92 $\pm$ 0.39 13.95 $\pm$ 4.09
Corvi23-R [Corvi23] 99.77 $\pm$ 0.40 3.67 $\pm$ 1.38
Corvi23-S [Corvi23] 9.12 $\pm$ 4.06 2.52 $\pm$ 1.47
Song24-Freq [Song24] 3.31 $\pm$ 1.38 6.66 $\pm$ 0.88
Qian20 [Qian20] 18.12 $\pm$ 15.79 12.51 $\pm$ 1.64
Wang20 [Wang20] 99.55 $\pm$ 0.77 5.68 $\pm$ 5.09
Song24-SL [Song24] 10.09 $\pm$ 1.90 7.29 $\pm$ 1.47
VI-B Robustness to Forgery
Table III shows that forgery ASRs are below those of removal across the board. Forgery is a fundamentally harder task than removal: with 12 models in the attribution pool, successful forgery requires targeted misattribution to a specific model, while removal only requires misclassification into any incorrect class. Geometrically, removal corresponds to crossing the nearest decision boundary, whereas forgery demands reaching a potentially narrow and distant target region. When only two models are considered (which is the case for general deepfake detection tasks), the two objectives coincide and removal corresponds to forgery, but as the number of models grows, the decision landscape becomes increasingly complex and harder to manipulate by a forgery attack.
Despite this difficulty, many methods still exhibit substantial forgery ASRs, especially in white-box settings. On average, W1 yields the highest ASRs, followed by W2 and then W3, consistent with their decreasing levels of gradient access—from direct gradients (W1) to analytic approximations (W2) to surrogate-based approximations (W3). Even in the black-box setting, ASRs reach up to 25.51%, indicating that a non-trivial fraction of samples can still be intentionally misattributed to a chosen target model. Importantly, note that the reported means and standard deviations are computed by averaging over all 12 generative models. Refer to Figure 4 for the variation in attack vulnerability of each model when used as a forgery target. In practice, the attacker may be interested in a specific model, or a specific model family that might be more vulnerable to attacks, in which case the average ASR within the family might be significantly higher. Furthermore, since our gradient-based attacks are generic and not tailored to the internal design of each method, method-specific adaptations could plausibly yield even higher forgery success rates.
Similar to removal attacks, forged samples cannot be visually distinguished from their original counterparts, maintaining an LPIPS <0.05<0.05 and PSNR >35>35 dB across all forged samples (see Tables VI, VII in Appendix A). It is important to note that higher ASR for forgery and removal could be achieved if the adversary relaxed our strict $\ell_{\infty}$ bound constraints, as explained in Appendix D.
Recall that we discarded B2 attack strategies from the forgery attack evaluation because they lack specificity to embed a targeted model fingerprint into the forged sample. For the remaining attacks, we observe trends similar to those in the removal evaluation. As with removal, white-box settings tend to outperform B1 strategies, with W1 consistently achieving the highest ASRs. In contrast, ASR variance for McCloskey18 is relatively lower than in removal attacks, while Qian20 now exhibits the highest variance. This indicates that while the objectives of removal and forgery are similar, they are not perfectly complementary—a finding that motivates a deeper investigation of their relationship.
VI-C Removal vs. Forgery
We observe a clear positive correlation between the ASR of removal and forgery, for both white-box and black-box scenarios. As shown in Fig. 2, MFD methods that are more susceptible to removal are also generally more vulnerable to forgery. This trend is intuitive: since all reported attacks are gradient-based, if gradients can efficiently guide perturbations toward a distant forgery region, they can likewise push samples outside their current decision region, enabling removal.
Figure 2: Correlation between removal and forgery attack vulnerabilities across MFD methods. Left: white-box scenario using the best ASR per method (highest ASR among W1, W2, and W3) for removal and forgery, capturing white-box vulnerability. Right: black-box I scenario (B1). Points are MFD methods; error bars show stdevs across runs. Forgery has no B2, so B2 is omitted. Colors encode fingerprint domain as shown in the legend box. See Figure 6 in Appendix B for individual whit-box attack data.
Figure 3: Trade-off between attribution utility (accuracy of the classifier $h$, y-axis) and robustness (ASR, x-axis) across four scenarios: removal under white-box (far left) and black-box I (second), and forgery under white-box (third) and black-box I (far right). White-box scenario uses the best ASR per method (highest ASR among W1, W2, and W3). Each point is an MFD method; error bars show standard deviation across runs. Methods nearer the top-left deliver stronger utility–robustness trade-offs. Green circles indicate outliers with better trade-off. See Figure 7 in Appendix C for individual whit-box attack data.
However, for the black-box scenario, this correlation does not imply that high removal ASR translates into high forgery ASR in absolute value terms. For B1 (Figure 2, Right), we observed that the increase in forgery ASR from less-resilient to more-resilient methods remains small in magnitude. The magnitude of this increase likely depends on the size of the model pool and the characteristics of the models themselves. For instance, in the simplest case of a two-model pool, the attribution problem reduces to binary classification (similar to deepfake detection). In this setting, removing fingerprints of one model effectively forces the sample to be attributed to the other, making removal and forgery effectively the same goal. By contrast, as the number of models in the pool grows, removal remains comparably feasible—since pushing samples outside the decision region of the true model does not fundamentally depend on the pool size and other models—while forgery can become increasingly difficult. This is because forging requires moving a sample into a specific target region, and the difficulty of reaching such regions grows with the number and arrangement of models’ decision boundaries.
Hence, both removal and forgery ASR scales can be affected by the geometric landscape of decision regions across models in each individual model attribution task. The precise relationship between removal and forgery thus depends on the size of the model pool and the characteristics of the models. A more detailed analysis of how pool size and model diversity govern this relationship is left to closer examination.
VI-D Utility–Robustness Trade-off
We benchmark the original attribution accuracy (clean, non-adversarial) of all evaluated MFD methods, with results summarized in Table I under the “Accuracy” column. A few approaches, such as [Corvi23, McCloskey18], were originally designed for binary deepfake detection (e.g., real vs. AI-generated images) rather than multi-class attribution. As a result, they offer limited utility for attribution tasks, despite the strong detection performance originally reported in the deepfake setting.
To examine the trade-off between attribution accuracy and adversarial robustness for MFD methods, we plot this relationship for both removal and forgery attacks in Figure 3 under white-box and black-box scenarios respectively. Broadly, we observe that the utility–robustness tension is pronounced for both removal and forgery attacks: methods with higher attribution accuracy tend to suffer higher ASR, indicating weaker robustness. Most high-utility methods (accuracy >0.7>0.7) exhibit significantly higher vulnerability than low-utility ones.
For example, neural network–based methods such as Wang20 achieve the highest attribution accuracy; yet, it is also among the most fragile methods to both removal and forgery, particularly under white-box attacks. Nonetheless, there are notable exceptions: for instance, Marra19a and Song24-RGB under black-box attacks, and Song24-Freq under white-box attacks. These methods pair moderate attribution accuracy with substantially stronger robustness, making them stand out as more balanced candidates among the evaluated approaches.
Overall, these findings highlight a fundamental trade-off: methods with high attribution accuracy often lack robustness, whereas more robust methods typically sacrifice accuracy. Promising directions for future work include: (i) improving residual- and manifold-based approaches such as Marra19a and Song24 to boost accuracy; (ii) strengthening CNN-based methods like Wang20 against adversarial perturbations; and (iii) exploring hybrid or ensemble designs that combine the robustness of residual/manifold methods with the strong attribution accuracy of CNN-based approaches.
VI-E Attack Variance Across Models
Figure 4: Model vulnerability per MFD method, averaged over all applicable attack types. Each subplot plots removal ASR (x) vs. forgery ASR (y) for the generative models (indicated by colors and markers) tied to that method; ASR spans 0–1 and error bars show standard deviation across different attack types. The final subplot averages the data over all MFD methods (with standard deviation representing statistics over different MFD methods), showing general vulnerability trend of generative models - models with more resilience to removal are more vulnerable to forgery, and vice versa.
To examine whether particular generative models are inherently more vulnerable to certain adversarial attacks, we evaluated per-model vulnerability across all MFD methods. Figure 4 plots forgery ASR versus removal ASR for each model under each method, averaging over all applicable attack strategies. We define a model’s removal vulnerability as how easily its own images can be perturbed to be misclassified as coming from any other model, and its forgery vulnerability as how easily images from other models can be perturbed to be misclassified as coming from this model.
The results show that vulnerability is not consistently high or low for any given model or model family. Instead, it varies strongly with the MFD method. For example, in removal, NVAE is among the least vulnerable under McCloskey18 and Song24-Freq, yet appears highly vulnerable under other methods. Likewise, ADM is markedly more robust in Nowroozi22 than all other models, but not in any of the remaining methods. In forgery, ADM is significantly more vulnerable under Nataraj19 and Nowroozi22, yet this pattern does not generalize across all MFD methods.
Despite this variability, an important and obvious trend emerges: across MFD methods, image generation models that are more robust to removal tend to be more vulnerable to forgery, and vice versa. This trade-off is visible in the final subplot of Figure 4, and is also particularly clear under methods such as Nowroozi22, Song24-Freq, and Qian20. Intuitively, if a model’s fingerprint occupies a more isolated decision region, it may be easy for gradients to escape that region (removal) but harder to steer perturbations into it (forgery). We further observe that diffusion and VAE models are more resistant to removal and more vulnerable to forgery, whereas GAN-based models show the opposite trend. We leave a deeper investigation of these behaviors to future work.
In summary, vulnerability patterns in attribution attacks cannot be attributed solely to the generative model architecture nor solely to the MFD method used for attribution. Instead, they arise from the interaction between the two, jointly shaping the attribution decision landscape.
VII Discussion
Researchers have developed MFD techniques to provide provenance of model outputs and thus address some of the accountability challenges posed by today’s increasingly powerful generative models. In this paper, we provide the first systematic security analysis of these techniques, which includes over 14 MFD techniques across a range of possible threat models and attack strategies. Despite the reported high accuracy of these techniques in a clean setting in the literature, our findings reveal that they are highly vulnerable to adversarial manipulations.
Our results show that most techniques fail to provide reliable model attribution under removal attacks. For most MFD techniques, our attacks remove model fingerprints across various access assumptions without visually degrading the attacked images. Most concerning, simple black-box image perturbations succeed in removing fingerprints, with attacks achieving over $90\%$ success for some techniques. These findings indicate that fingerprint traces can be disrupted without sophisticated attacks or knowledge of the underlying MFD system.
Implications for MFD Technique Deployment. These results fundamentally undermine the applications proposed in the academic literature. The stark contrast between MFD’s high accuracy in the clean setting (often $90\%$) and its fragility under attack, uncovers a gap that can lead to a false sense of security if unacknowledged. For example, such high attack success rates imply that investigators and model providers cannot use these techniques to reliably trace images to the source models, compromising their forensic value and undermining the accountability they are intended to provide. Importantly, while forgery attacks are significantly less successful than removal, effective removal alone severely compromises MFD’s value for such applications.
Even in the clean setting, existing evaluations may have overestimated MFD performance for model attribution. Current studies evaluate small closed worlds of at most 20 different models, raising questions about how representative they are of real-world scenarios. Practical deployments could be far more complex. For example, detecting the source of an arbitrary image would require distinguishing among orders of magnitude more models. Furthermore, real-world scenarios might involve highly similar models: the ML community often reuses public datasets for training and relies on open sourced learning algorithms, potentially producing similar model fingerprints—a situation not represented by the small, diverse model pools considered in the literature. Conversely, other practical scenarios might involve smaller model pools. Additional sources of evidence can narrow down the candidate pool to one closer to those in the literature, potentially enabling effective MFD deployment. Therefore, we encourages future papers presenting new MFD schemes to evaluate them in realistic open worlds. However, the success of our attacks implies that these techniques remain unreliable even in such ideal conditions, highlighting the urgent need for developing MFD techniques that are robust in adversarial environments.
Implications for the Adversarial Arms-Race. Our results also have implications for the strategies chosen by attackers and fingerprinters. Since attackers cannot predict which MFD technique will be applied, they must develop attacks capable of removing fingerprints across all MFD techniques. This situation might seem like a whack-a-mole game that favors the fingerprinter, who can apply multiple techniques simultaneously. However, our results add nuances to this: we observe a tension between utility and robustness, where the most robust methods (e.g., Marra19a, Song24-RGB) exhibit higher detection error than the more vulnerable ones. Consequently, even when the fingerprinter deploys all 14 techniques, the attacks increase overall attribution uncertainty and thus undermine value of these techniques for forensic applications.
Limitations and Future Work. One of the main limitations of our work is that we cannot conclude that the removal attacks completely erase the fingerprints: they supress enough signal to impair model attribution under a specific experimental setting, but there may be enough signal to still enable deepfake detection. Future work could further investigate the effectiveness of the attacks against MFD techniques for deepfake detection.
Another avenue for future work is a more thorough exploration of the attack hyperparameter space. Although we tuned parameters that had significant impact, we made no attempt at optimization. Similarly, although future work could explore attacks tailored to specific techniques, attackers must be strategic: since attackers cannot predict which MFD techniques fingerprinters will deploy, attackers are better off increasing overall maximum success across techniques. Therefore, even highly successful bespoke attacks provide only marginal advantage if other techniques remain robust or if increased success against one method comes at the cost of reduced effectiveness against others.
From the perspective of mitigations, an important direction for future work is to evaluate defenses from adversarial machine learning, such as adversarial training or randomized smoothing, applied to the attribution model. A thorough investigation would require adapting these defenses to handle both removal and forgery attacks and developing corresponding adaptive attacks to properly assess robustness. Such methods must also consider the typical utility–robustness trade-off, as improved robustness often comes at the cost of reduced clean accuracy. We view this as a substantial research effort that merits a dedicated study, and thus leave the systematic design and evaluation of defenses for future work.
Additionally, we suspect that fingerprint suppression may introduce detectable patterns. We leave for future work investigating the detectability of the attacks. This capability would be particularly valuable for countering forgery attacks, as early detection could prevent adversaries from successfully misattributing content. Another avenue for defense is to explore hybrid approaches. In particular, despite their lower detection accuracies, our results identify Marra19a and Song24-RGB’s fingerprint features as particularly robust under black-box conditions. Future work could refine the features to maintain robustness while improving performance.
Active vs. Passive Provenance Mechanisms. Watermarking has emerged as an alternative to fingerprinting for providing similar accountability guarantees in the deployment of generative models. The increased control of watermarking over the identifying signal makes active techniques more suitable for adversarial environments. Recently proposed watermarking schemes demonstrate robustness against sophisticated removal attempts while providing undetectability and formal security guarantees [yang2024gaussian, gunn2024undetectable]. On the other hand, watermarking requires integration during model training or inference, which might introduce friction in its adoption by the industry. We encourage future work to further explore trade-offs between these two types of techniques or strengthen performance and robustness by combining the best of both worlds.
VIII Conclusion
Model fingerprint detection (MFD) has been proposed as a means of providing provenance and accountability in the deployment of image generation models, with reported accuracies often exceeding $90\%$ in non-adversarial evaluations. However, our findings show that such results offer an incomplete picture of their reliability in the adversarial settings where attribution mechanisms are most critical. Through a systematic evaluation of 14 MFD techniques against 5 attack strategies, we expose a critical security gap: adversaries can reliably suppress or forge fingerprints, undermining the practical utility of current MFD methods for real-world attribution. While industry adoption of attribution mechanisms is already underway, it is unlikely to pause until robust solutions emerge. This urgency underscores the need not only to advance more resilient MFD techniques but also to develop complementary or alternative accountability mechanisms that remain effective under adversarial conditions. Looking forward, we identify several promising directions for strengthening model provenance of AI-generated images, including but not limited to leveraging more inherently robust methods such as residual- or manifold-based approaches, adversarially robust training of MFD systems, and hybrid designs that combine complementary strengths across existing model attribution techniques such as model watermarking.
Usage of LLMs
LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality.
Acknowledgments
This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh. Access to EIDF was facilitated through the University of Edinburgh’s Generative AI Laboratory GAIL Fellow scheme.
Appendix A Image Quality Under Attacks
TABLE IV: Perceptual distances (LPIPS, mean $\pm$ std) of removal attacks against evaluated MFD methods. White-box attacks include direct gradient-based optimization (W1), analytic approximations (W2), and surrogate fingerprint extractors (W3). Black-box attacks include surrogate classifiers (B1) and fingerprint-agnostic perturbations (B2: Gaussian noise, blur, JPEG compression, resizing). Lower LPIPS indicates better perceptual quality preservation.
Method W1 W2 W3 B1 B2-Noise B2-Blur B2-JPEG B2-Resize
McCloskey18 [McCloskey18] 0.0261 $\pm$ 0.0011 0.0247 $\pm$ 0.0016 0.0270 $\pm$ 0.0012 0.0246 $\pm$ 0.0007 0.0176 $\pm$ 0.0006 0.0286 $\pm$ 0.0006 0.0175 $\pm$ 0.0004
Nataraj19 [Nataraj19] 0.0238 $\pm$ 0.0010 0.0242 $\pm$ 0.0013 0.0263 $\pm$ 0.0011 0.0147 $\pm$ 0.0003 0.0247 $\pm$ 0.0005 0.0320 $\pm$ 0.0006 0.0262 $\pm$ 0.0002
Nowroozi22 [Nowroozi22] 0.0251 $\pm$ 0.0019 0.0246 $\pm$ 0.0015 0.0285 $\pm$ 0.0020 0.0153 $\pm$ 0.0003 0.0242 $\pm$ 0.0005 0.0312 $\pm$ 0.0005 0.0252 $\pm$ 0.0005
Song24-RGB [Song24] 0.0239 $\pm$ 0.0005 0.0275 $\pm$ 0.0016 0.0146 $\pm$ 0.0003 0.0248 $\pm$ 0.0004 0.0320 $\pm$ 0.0005 0.0263 $\pm$ 0.0007
Marra19a [Marra19a] 0.0256 $\pm$ 0.0010 0.0263 $\pm$ 0.0014 0.0271 $\pm$ 0.0009 0.0147 $\pm$ 0.0004 0.0249 $\pm$ 0.0004 0.0321 $\pm$ 0.0002 0.0260 $\pm$ 0.0003
Durall20 [Durall20] 0.0268 $\pm$ 0.0020 0.0275 $\pm$ 0.0018 0.0259 $\pm$ 0.0012 0.0143 $\pm$ 0.0005 0.0255 $\pm$ 0.0005 0.0322 $\pm$ 0.0007 0.0271 $\pm$ 0.0003
Dzanic20 [Dzanic20] 0.0227 $\pm$ 0.0011 0.0269 $\pm$ 0.0014 0.0254 $\pm$ 0.0010 0.0153 $\pm$ 0.0004 0.0246 $\pm$ 0.0005 0.0316 $\pm$ 0.0004 0.0253 $\pm$ 0.0003
Giudice21 [Giudice21] 0.0272 $\pm$ 0.0010 0.0265 $\pm$ 0.0008 0.0147 $\pm$ 0.0004 0.0248 $\pm$ 0.0006 0.0322 $\pm$ 0.0004 0.0263 $\pm$ 0.0004
Corvi23-R [Corvi23] 0.0281 $\pm$ 0.0009 0.0258 $\pm$ 0.0007 0.0168 $\pm$ 0.0003 0.0243 $\pm$ 0.0006 0.0323 $\pm$ 0.0005 0.0250 $\pm$ 0.0004
Corvi23-S [Corvi23] 0.0265 $\pm$ 0.0025 0.0249 $\pm$ 0.0020 0.0263 $\pm$ 0.0005 0.0206 $\pm$ 0.0008 0.0354 $\pm$ 0.0004 0.0220 $\pm$ 0.0004
Song24-Freq [Song24] 0.0236 $\pm$ 0.0011 0.0267 $\pm$ 0.0006 0.0146 $\pm$ 0.0002 0.0247 $\pm$ 0.0005 0.0319 $\pm$ 0.0006 0.0257 $\pm$ 0.0007
Qian20 [Qian20] 0.0284 $\pm$ 0.0003 0.0269 $\pm$ 0.0010 0.0141 $\pm$ 0.0001 0.0254 $\pm$ 0.0002 0.0313 $\pm$ 0.0001 0.0260 $\pm$ 0.0003
Wang20 [Wang20] 0.0278 $\pm$ 0.0002 0.0272 $\pm$ 0.0068 0.0142 $\pm$ 0.0003 0.0254 $\pm$ 0.0001 0.0310 $\pm$ 0.0005 0.0264 $\pm$ 0.0002
Song24-SL [Song24] 0.0241 $\pm$ 0.0050 0.0274 $\pm$ 0.0043 0.0145 $\pm$ 0.0004 0.0248 $\pm$ 0.0003 0.0316 $\pm$ 0.0002 0.0257 $\pm$ 0.0003
TABLE V: Peak signal-to-noise ratio (PSNR, mean $\pm$ std, in dB) of removal attacks against evaluated MFD methods. White-box attacks include direct gradient-based optimization (W1), analytic approximations (W2), and surrogate fingerprint extractors (W3). Black-box attacks include surrogate classifiers (B1) and fingerprint-agnostic perturbations (B2: Gaussian noise, blur, JPEG compression, resizing). Higher PSNR indicates better perceptual quality preservation.
Method W1 W2 W3 B1 B2-Noise B2-Blur B2-JPEG B2-Resize
McCloskey18 [McCloskey18] 37.4012 $\pm$ 0.0674 36.2441 $\pm$ 0.5015 36.7965 $\pm$ 0.0203 42.0924 $\pm$ 0.0251 40.2988 $\pm$ 0.1299 37.5891 $\pm$ 0.0749 40.0573 $\pm$ 0.1292
Nataraj19 [Nataraj19] 35.6142 $\pm$ 0.0021 36.9385 $\pm$ 0.1912 35.9843 $\pm$ 0.0593 41.8021 $\pm$ 0.0061 40.5872 $\pm$ 0.0541 37.1992 $\pm$ 0.0330 40.1264 $\pm$ 0.0562
Nowroozi22 [Nowroozi22] 35.8034 $\pm$ 0.8610 36.9541 $\pm$ 0.2201 35.9012 $\pm$ 0.6812 41.8362 $\pm$ 0.0171 40.6510 $\pm$ 0.1518 37.2914 $\pm$ 0.0903 40.2135 $\pm$ 0.1486
Song24-RGB [Song24] 37.0224 $\pm$ 0.0049 36.8610 $\pm$ 2.0319 41.8074 $\pm$ 0.0091 40.5244 $\pm$ 0.0443 37.1814 $\pm$ 0.0201 40.0691 $\pm$ 0.0369
Marra19a [Marra19a] 35.1841 $\pm$ 0.0571 35.5762 $\pm$ 0.6992 35.6635 $\pm$ 0.0078 41.7992 $\pm$ 0.0061 40.5863 $\pm$ 0.0694 37.2075 $\pm$ 0.0294 40.1234 $\pm$ 0.0619
Durall20 [Durall20] 35.2911 $\pm$ 0.6684 35.4324 $\pm$ 0.6834 35.4681 $\pm$ 0.2861 41.7613 $\pm$ 0.0198 40.6843 $\pm$ 0.0889 37.2412 $\pm$ 0.0626 40.2044 $\pm$ 0.0953
Dzanic20 [Dzanic20] 35.3134 $\pm$ 1.7921 35.2542 $\pm$ 0.0094 36.8073 $\pm$ 0.0165 41.8053 $\pm$ 0.0095 40.6964 $\pm$ 0.0778 37.3095 $\pm$ 0.0386 40.2531 $\pm$ 0.0710
Giudice21 [Giudice21] 35.2413 $\pm$ 0.0061 35.2544 $\pm$ 0.0096 41.7913 $\pm$ 0.0063 40.5821 $\pm$ 0.0561 37.1963 $\pm$ 0.0310 40.1192 $\pm$ 0.0549
Corvi23-R [Corvi23] 35.2704 $\pm$ 0.0171 35.5209 $\pm$ 0.0579 41.8264 $\pm$ 0.0187 40.6084 $\pm$ 0.0891 37.2962 $\pm$ 0.0478 40.1794 $\pm$ 0.0809
Corvi23-S [Corvi23] 35.3019 $\pm$ 0.0304 35.4150 $\pm$ 0.0094 41.8374 $\pm$ 0.0224 41.8022 $\pm$ 0.2835 38.1401 $\pm$ 0.1474 41.4723 $\pm$ 0.2899
Song24-Freq [Song24] 36.9454 $\pm$ 0.1692 35.4734 $\pm$ 0.0063 41.8013 $\pm$ 0.0060 40.5971 $\pm$ 0.0561 37.2188 $\pm$ 0.0317 40.1374 $\pm$ 0.0548
Qian20 [Qian20] 35.2674 $\pm$ 0.0012 35.4744 $\pm$ 0.4161 41.8034 $\pm$ 0.0060 40.6004 $\pm$ 0.0012 37.1884 $\pm$ 0.0011 40.1184 $\pm$ 0.0009
Wang20 [Wang20] 35.2685 $\pm$ 0.0013 35.5591 $\pm$ 0.4302 41.8103 $\pm$ 0.0011 40.6022 $\pm$ 0.0009 37.1871 $\pm$ 0.0008 40.1181 $\pm$ 0.0008
Song24-SL [Song24] 35.8811 $\pm$ 0.8534 35.4968 $\pm$ 0.1459 41.8101 $\pm$ 0.0090 40.5662 $\pm$ 0.0604 37.2110 $\pm$ 0.0313 40.1133 $\pm$ 0.0580
TABLE VI: Perceptual distances (LPIPS, mean $\pm$ std) of forgery attacks against evaluated MFD methods. White-box attacks include direct gradient-based optimization (W1), analytic approximations (W2), and surrogate fingerprint extractors (W3). Black-box attacks include surrogate classifiers (B1). Lower LPIPS indicates better perceptual quality preservation.
Method W1 W2 W3 B1
McCloskey18 [McCloskey18] 0.0264 $\pm$ 0.0011 0.0258 $\pm$ 0.0010 0.0271 $\pm$ 0.0012
Nataraj19 [Nataraj19] 0.0243 $\pm$ 0.0003 0.0238 $\pm$ 0.0003 0.0246 $\pm$ 0.0003
Nowroozi22 [Nowroozi22] 0.0252 $\pm$ 0.0012 0.0247 $\pm$ 0.0013 0.0256 $\pm$ 0.0013
Song24-RGB [Song24] 0.0241 $\pm$ 0.0007 0.0228 $\pm$ 0.0008
Marra19a [Marra19a] 0.0257 $\pm$ 0.0009 0.0249 $\pm$ 0.0008 0.0244 $\pm$ 0.0006
Durall20 [Durall20] 0.0246 $\pm$ 0.0030 0.0239 $\pm$ 0.0018 0.0232 $\pm$ 0.0029
Dzanic20 [Dzanic20] 0.0250 $\pm$ 0.0010 0.0263 $\pm$ 0.0012 0.0257 $\pm$ 0.0011
Giudice21 [Giudice21] 0.0268 $\pm$ 0.0009 0.0246 $\pm$ 0.0007
Corvi23-R [Corvi23] 0.0284 $\pm$ 0.0012 0.0261 $\pm$ 0.0014
Corvi23-S [Corvi23] 0.0269 $\pm$ 0.0020 0.0279 $\pm$ 0.0009
Song24-Freq [Song24] 0.0240 $\pm$ 0.0007 0.0243 $\pm$ 0.0009
Qian20 [Qian20] 0.0265 $\pm$ 0.0002 0.0244 $\pm$ 0.0002
Wang20 [Wang20] 0.0271 $\pm$ 0.0002 0.0240 $\pm$ 0.0002
Song24-SL [Song24] 0.0251 $\pm$ 0.0010 0.0264 $\pm$ 0.0102
TABLE VII: Peak signal-to-noise ratio (PSNR, mean $\pm$ std, in dB) of forgery attacks against evaluated MFD methods. White-box attacks include direct gradient-based optimization (W1), analytic approximations (W2), and surrogate fingerprint extractors (W3). Black-box attacks include surrogate classifiers (B1). Higher PSNR indicates better perceptual quality preservation.
Method W1 W2 W3 B1
McCloskey18 [McCloskey18] 37.1634 $\pm$ 0.0439 37.0112 $\pm$ 0.0265 36.9873 $\pm$ 0.0224
Nataraj19 [Nataraj19] 36.7484 $\pm$ 0.0011 37.0125 $\pm$ 0.0093 36.7191 $\pm$ 0.0050
Nowroozi22 [Nowroozi22] 36.8803 $\pm$ 0.0268 37.0182 $\pm$ 0.0204 36.7652 $\pm$ 0.0156
Song24-RGB [Song24] 37.0231 $\pm$ 0.0058 37.2444 $\pm$ 0.0342
Marra19a [Marra19a] 36.7693 $\pm$ 0.0199 36.7671 $\pm$ 0.0187 36.7175 $\pm$ 0.0122
Durall20 [Durall20] 36.8368 $\pm$ 0.0539 36.7484 $\pm$ 0.0273 36.7133 $\pm$ 0.0548
Dzanic20 [Dzanic20] 36.3375 $\pm$ 0.0248 36.7914 $\pm$ 0.0200 36.7445 $\pm$ 0.0075
Giudice21 [Giudice21] 36.8524 $\pm$ 0.0089 36.7091 $\pm$ 0.0090
Corvi23-R [Corvi23] 36.7491 $\pm$ 0.0090 36.7804 $\pm$ 0.0092
Corvi23-S [Corvi23] 35.9128 $\pm$ 0.0638 36.7703 $\pm$ 0.0141
Song24-Freq [Song24] 37.0042 $\pm$ 0.0075 36.7284 $\pm$ 0.0089
Qian20 [Qian20] 36.7428 $\pm$ 0.0048 36.6938 $\pm$ 0.0090
Wang20 [Wang20] 36.7263 $\pm$ 0.0061 36.7602 $\pm$ 0.0051
Song24-SL [Song24] 36.7611 $\pm$ 0.0659 36.7183 $\pm$ 0.1011
LPIPS and PSNR indicators. To complement the evaluation of attack performance, we also assess the visual quality of adversarially perturbed images. We report LPIPS, a perceptual similarity metric, and PSNR, a pixel-level fidelity measure, as two widely adopted indicators of image quality under perturbations. Tables IV and V summarize the results for removal attacks, while Tables VI and VII present the corresponding results for forgery attacks. Across both attack goals and all attack types and MFD methods, LPIPS values remain below 0.05 and PSNR values consistently exceed 35 dB, indicating that adversarial perturbations preserve high image fidelity.
Example Images. As an example, Figure 5 provides a comprehensive visualization of successful removal attacks against the Song24-Freq MFD method across different generative models and attack strategies. The grid structure clearly illustrates how various attack types (W3, B1, and B2 with different perturbation techniques) perform on different target models, with each cell showing the original image, the perturbed result, and the scaled difference map. This visualization demonstrates the effectiveness of different attack strategies in removing model fingerprints while preserving the visual quality of the generated images. The difference images, scaled for visibility, reveal the spatial distribution of perturbations introduced by each attack method.
Figure 5: Visualization of successful removal attacks against the Song24-Freq MFD method for selected models. Each row represents a different generative model (target for fingerprint removal), and each column represents a different attack type. Within each cell, the three images show (from left to right): the original image, the perturbed image after attack, and the scaled difference. The difference is linearly mapped such that the lowest value corresponds to black and the highest value corresponds to white, with intermediate values scaled accordingly. The attack types include W3 (surrogate extractor), B1 (surrogate classifier), and B2 (basic image perturbations: blur, Gaussian noise, JPEG compression, and resize). The visualization demonstrates the effectiveness of various attack strategies in removing model fingerprints while maintaining visual quality. An “N/A” entry indicates non-existence of a successfully attacked image for the case.
Appendix B Removal ASR vs Forgery ASR
Here in Figure 6 we report the relationship between removal-attack ASR and forgery-attack ASR for each gradient-based attack type separately, rather than aggregating them using the best white-box ASR as done in the main text (i.e. Figure 2). Specifically, we present results for W1, W2, and W3 individually. Note that certain attack types are only applicable to specific fingerprinting methods (e.g., W1 requires differentiability and is therefore only compatible with MFD methods such as Giudice21, Corvi23-R, Corvi23-S, Qian20, and Wang20). Importantly, the strong positive correlation between removal and forgery ASR persists for each attack type considered independently, reinforcing the robustness of the trend reported in the main paper.
Figure 6: Correlation between removal and forgery attack vulnerabilities across MFD methods. We plot the data with regard to each attack type separately in this figure, as opposed to Figure 2. Points are MFD methods; error bars show standard deviation across 5 independent runs. Forgery has no B2, so B2 is omitted. Colors encode fingerprint domain: blue=RGB, orange=frequency, purple=NN.
Appendix C Utility-Robustness Trade-off
Here in Figure 7 we report the trade-off relationship between utility and attack ASR for each gradient-based attack type separately, rather than aggregating them using the best white-box ASR as done in the main text (i.e. Figure 3). Specifically, we present results for W1, W2, and W3 individually. Note that certain attack types are only applicable to specific fingerprinting methods.
Figure 7: Trade-off between attribution utility (accuracy of the classifier $h$, y-axis) and robustness (ASR, x-axis). We plot the data with regard to each attack type separately in this figure, as opposed to Figure 3. Each point is an MFD method; error bars show standard deviation across 5 independent runs. Methods nearer the top-left deliver stronger utility–robustness trade-offs.
Appendix D Hyperparameter Tuning
We conducted hyperparameter tuning for several attack cases with respect to their strength parameters to illustrate the trade-off between ASR and image quality degradation. Specifically, we report ablation results for multiple attack cases using MFD Dzanic20 as a representative example, as shown in Figures 8 and 9.
For PGD-based attacks (W3 and B1), we varied the perturbation budget $\epsilon$ (the maximum allowed change per pixel) over the range {0.001, 0.005, 0.01, 0.05, 0.1, 0.5, 1.0}.
For image-perturbation (B2) attacks, we swept each perturbation parameter from mild to strong settings:
  • Gaussian noise standard deviation: {0.001, 0.0025, 0.005, 0.01, 0.02};
  • Gaussian blur sigma: {0.2, 0.5, 1.0, 1.5};
  • JPEG quality: {95, 90, 85, 80, 75, 70, 65, 50};
  • Resize scaling factor: {0.9, 0.8, 0.7, 0.6, 0.5}.
Figures 8 and 9 plot ASR and image quality indicators (LPIPS/PSNR) as functions of these attack-strength hyperparameters. As expected, increasing attack strength generally yields higher ASR, but at the cost of noticeably degraded image quality, particularly for PGD-based attacks.
[…]

### v2

@@
 Abstract
-Model fingerprint detection techniques have emerged as a promising approach for attributing AI-generated images to their source models, but their robustness under adversarial conditions remains largely unexplored.
+Model fingerprint detection has shown promise to trace the provenance of AI-generated images in forensic applications.
+However, despite the inherent adversarial nature of these applications, existing evaluations rarely consider adversarial settings.
 We present the first systematic security evaluation of these techniques, formalizing threat models that encompass both white- and black-box access and two attack goals: fingerprint removal, which erases identifying traces to evade attribution, and fingerprint forgery, which seeks to cause misattribution to a target model.
@@
 Our experiments reveal a pronounced gap between clean and adversarial performance.
-Removal attacks are highly effective, often achieving success rates above $80\%$ in white-box settings and over $50\%$ under constrained black-box access.
-While forgery is more challenging than removal, its success significantly varies across targeted models.
-We also identify a utility–robustness trade-off: methods with the highest attribution accuracy are often vulnerable to attacks.
-Although some techniques exhibit robustness in specific settings, none achieves high robustness and accuracy across all evaluated threat models.
-These findings highlight the need for techniques balancing robustness and accuracy, and identify the most promising approaches for advancing this goal.
+Removal attacks are highly effective, often achieving success rates above $80\%$ in white-box settings and over $50\%$ under black-box access.
+While forgery is more challenging than removal, its success varies significantly across targeted models.
+We also observe a utility–robustness trade-off: accurate attribution methods are often vulnerable to attacks and, although some techniques are robust in specific settings, none achieves robustness and accuracy across all evaluated threat models.
+These findings highlight the need for techniques that balance robustness and accuracy, and we identify the most promising approaches toward this goal.
+Code available at: https://github.com/kaikaiyao/SmudgedFingerprints.
 I Introduction
@@
 While these technological advances enable new forms of creativity, they also come with risks for misinformation, fraud, and abuse, particularly when AI-generated content is passed as genuine [bird2023typology].
-The challenge of reliably identifying and tracing the generated images has thus become critical for maintaining trust and accountability in digital media.
+As a consequence, reliably identifying and tracing generated images has become a pressing challenge for maintaining trust and accountability in digital media.
 A promising avenue to address this challenge builds on a key property of these models: they tend to leave artifacts—subtle but consistent patterns—in the images they produce [Marra18].
@@
 Existing MFD techniques evaluated under various settings achieve over 90% accuracy for both deepfake detection [Marra19a, Nataraj19, Durall20, Dzanic20, Wang20, Corvi23] and model attribution [Marra18, Marra19b, Yu19, Wang20, Guarnera20, Girish21].
-If these accuracies were to hold in realistic scenarios, they would prove useful in bridging the accountability gap in current deployments of generative models.
-Indeed, being able to identify deepfakes would help mitigate the risks of misinformation and scams, while attributing a harmful output to the specific model that generated it would make it possible to flag issues, understand their causes, and prevent similar problems in the future.
+If these accuracies were to hold in realistic scenarios, MFD would prove useful in bridging the accountability gap in current deployments of generative models.
+Indeed, being able to identify deepfakes would help mitigate the risks of misinformation and scams, while attributing a harmful output to the specific model that generated it would make it possible to flag issues, understand their causes, and prevent similar issues in the future.
 However, high detection accuracy alone is not sufficient to guarantee the practical utility of these techniques.
@@
 These threat models lead to five attack strategies that we implement and evaluate on a comprehensive set of MFD techniques compiled from the literature, spanning different feature extraction approaches (pixel-domain, frequency-domain, learned representations), detection algorithms (statistical methods, neural networks), and evaluation settings.
-We focus on model attribution rather than deepfake detection because it is a harder problem: model attribution does not merely require detecting the presence of model-induced patterns, but to detect architecture-specific differences across these patterns.
-Therefore, model attribution is at least as hard as deepfake detection and by evaluating MFD techniques in this setting, we evaluate the worst-case scenario for the fingerprinter, allowing us to draw conclusions for the other scenario: if the techniques prove robust against attacks on model attribution, they should remain robust for the simpler task of deepfake detection.
+We focus on model attribution rather than deepfake detection because model attribution is the harder task for MFD: it requires not merely detecting the presence of model-induced patterns, but distinguishing architecture-specific differences across them, which requires higher entropy.
+Therefore, evaluating MFD techniques for model attribution allows us to draw conclusions for deepfake detection: if the techniques prove robust against attacks on model attribution, they should remain robust in detecting for the simpler task of deepfake detection.
 Our threat models consider two attacker goals: fingerprint removal and fingerprint forgery—the two attacks that have been evaluated in the literature on watermarking for deepfake detection.
@@
 Unlike watermarking, which embeds signals during generation, MFD relies on detecting inherent artifacts or “fingerprints” left by the model’s processes.
-These fingerprints can be detected in statistical irregularities of color and frequency patterns, as well as in feature representations.
+These fingerprints can be detected in statistical irregularities of color and frequency patterns in the image.
 Following the classification in Song et al.
@@
 For example, approaches using azimuthally averaged power spectral densities from 2D FFTs have demonstrated near-perfect discrimination (close to 100% accuracy) for deepfake detection between GAN outputs and real images on CelebA and LSUN, detecting patterns in high-frequency components for models like DCGAN and CycleGAN [Durall20].
-Other methods extract features from Fourier modes, such as decay exponents and magnitudes, achieving high accuracy (up to 99%) for deepfake detection on ProGAN-generated faces [Dzanic20].
+Other methods extract features from Fourier modes, such as decay exponents and magnitudes, achieving up to 99% accuracy for deepfake detection on ProGAN-generated faces [Dzanic20].
 Giudice et al.
@@
 While MFD methods perform well in benign settings, their robustness against adversarial evasion remains underexplored—particularly for model attribution, where attackers may attempt to remove or forge subtle, model-specific fingerprints.
-Existing evaluations largely examine resilience to generic image perturbations or targeted fingerprint removal, typically in black-box settings that do not adapt to the detector’s architecture.
-Moreover, most studies emphasize deepfake detection rather than attribution.
+Existing evaluations mostly test robustness against generic image perturbations or targeted fingerprint removal, typically in black-box settings that do not adapt to the detector’s architecture.
+Moreover, most studies emphasize deepfake detection rather than model attribution.
 Regarding the study of attacks, prior work has almost exclusively focused on fingerprint removal, overlooking forgery attacks in which adversaries mimic legitimate fingerprints to mislead attribution.
 Generic Image Perturbations.
-Most evaluations test MFD robustness against generic image perturbations such as JPEG compression, blurring, resizing, and rotation, which can unintentionally degrade fingerprints.
+Most evaluations evalate MFD robustness to generic image perturbations such as JPEG compression, blurring, resizing, and rotation, which can unintentionally degrade fingerprints.
 In the spatial pixel domain, Fourier spectrum–based methods have shown that JPEG compression can reduce detection accuracy by 10–20% on ProGAN datasets, primarily due to the loss of high-frequency components [Dzanic20].
@@
 Method-Specific Attacks.
-A limited number of studies have investigated adaptive attacks specifically designed against MFD techniques, aiming to suppress fingerprints while maintaining perceptual quality.
+A few studies have investigated adaptive attacks specifically designed against MFD techniques, aiming to suppress fingerprints while maintaining perceptual quality.
 For co-occurrence-based detectors such as Nataraj et al.’s [Nataraj19], projected gradient descent (PGD) approximations of the co-occurrence operator can adjust pixel statistics to resemble those of real images, reducing detection accuracy from over 99% to nearly 0% in white-box settings on CycleGAN datasets [Goebel20].
@@
 These results demonstrate the vulnerability of MFD methods to adaptive adversaries.
-However, the extent of their susceptibility to adaptive attacks, particularly against forgery and within attribution settings, remains largely underexplored.
-II-C Research Gap
+However, the extent of their susceptibility to adaptive attacks, particularly against forgery and within model attribution settings, remains largely underexplored.
+II-C Research Gaps
 Most MFD studies overlook the security of the techniques despite that the authors propose inherently adversarial applications.
@@
     All prior research has focused on removal attacks, leaving robustness to forgery attacks underexplored.
-As we explain in Section III-B, although these two attacks are related, effectiveness of one does not imply the other.
+As we explain in Section III-B, although these two attacks are related, success of one does not imply the other.
 Our work closes this gap by formalizing threat models for MFD techniques and systematically deriving attack strategies from these threat models.
@@
 Some approaches instead employ end-to-end neural networks, where fingerprint extraction is not a separate function (e.g., [Qian20, Wang20]).
-In such cases, the fingerprint can be defined implicitly: the preceding layers act as the fingerprint extractor $\phi$, while the final classification layer serves as the model attribution classifier $h$.
+In such cases, the fingerprint can be defined implicitly: the neural network except the last layer acts as the fingerprint extractor $\phi$, and the last layer can be interpreted as the model attribution classifier $h$.
 III-B Adversarial Objectives
@@
 where $d:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{\geq 0}$ is a distance function and $\varepsilon$ specifies the perturbation budget.
-Typical choices of $d$ include $\ell_{\infty}$ or $\ell_{2}$ norms, LPIPS [zhang2018unreasonable], or PSNR.
-This constraint ensures that adversarial manipulations do not degrade the quality of the attacks images excessively.
+Typical choices of $d$ include $\ell_{\infty}$ and $\ell_{2}$ norms, LPIPS [zhang2018unreasonable], and PSNR.
+This constraint ensures that adversarial manipulations do not degrade the quality of the attacked images excessively.
 We consider two adversarial objectives under this constraint:
@@
     $\mathcal{F}(x^{\prime})\neq y,\quad d(x,x^{\prime})\leq\varepsilon.$ (3)
-    A successful removal attack suppresses enough model-specific information in the fingerprint features to cause model attribution classifier $h$ to misclassify $x^{\prime}$.
+    A successful removal attack suppresses enough model-specific information in the fingerprint features to cause the model attribution classifier $h$ to misclassify $x^{\prime}$.
 Since the MFD system operates in a closed world of candidate generators, such misclassification typically results in a low-confidence attribution to a different model.
@@
 In addition, although access to a generator might give the adversary an advantage against an MFD method designed for that specific generator, most MFD methods—like our attacks—are generator-agnostic and applicable across generator families.
-Note that removal attacks are feasible under all three knowledge levels, but forgery attacks are infeasible in the Black-box II setting because the adversary lacks access to $G_{t}$ and thus lacks the information needed to cause a targeted misattribution.
+Removal attacks are feasible under all three knowledge levels, while forgery attacks are infeasible in the Black-box II setting, as the adversary does not have access to $G_{t}$, and thus lacks the information needed to cause a targeted misattribution.
 IV Attack Implementation
@@
 Attacks then optimize $h(\phi_{s}(x))$.
-Figure 1 depicts which white-box attacks (W1–W3) apply given $\phi$’s properties.
-These strategies differ in their assumptions: W1 makes the strictest, requiring perfect gradient information.
+Figure 1 depicts which white-box attacks (W1–W3) apply given properties of $\phi$.
+These strategies differ in their assumptions: W1 is the strictest, requiring perfect gradient information.
 W2 and W3 are possible when there is gradient information but offer no advantage over W1.
@@
 Under Black-box Access II, without knowledge of $\phi$, $h$, or $\mathcal{Y}$, a typical adversary applies generic image transformations such as JPEG compression, noising, blurring, or resizing.
-These untargeted manipulations aim to degrade fragile fingerprint traces and reduce attribution confidence and without access to the target generator they cannot achieve forgery.
+These untargeted manipulations aim to degrade fragile fingerprint traces and reduce attribution confidence.
 Under Black-box Access II, adversaries could pursue more sophisticated attacks, such as training a surrogate on outputs from the target generator and a dataset for the negative class.
@@
   • B1: $\mathcal{F}^{\prime}(x)=h_{s}(x)$, where $h_{s}$ is a surrogate attribution classifier trained on images.
-The adversarial loss $\mathcal{L}_{\text{adv}}$ is defined according attack objective:
+The definition of the adversarial loss $\mathcal{L}_{\text{adv}}$ depends on the attack objective:
   • Fingerprint Removal: suppress attribution to the true source $y$ by optimizing a loss that discourages high confidence on that label. We use a cross-entropy–based objective:
@@
 We evaluate 14 MFD methods spanning the RGB, frequency, and learned feature domains.
-Although this set is not exhaustive, it is representative: it includes influential methods that have been widely studied for model attribution (e.g., those covered in Song et al.
+Although this set is not exhaustive, it is representative: it includes influential MFD methods that have been widely studied for model attribution (e.g., those covered in Song et al.
 [Song24]) and thus reflects the performance and robustness trends of the state of the art.
 Table I summarizes their key characteristics, including the feature domain (RGB, frequency, or learned representations [Song24]), a brief description of the fingerprint features, and the fingerprint size (i.e., the output of $\phi$, or - for end-to-end methods).
-We also report attribution accuracy of the MFD methods, obtained by reproducing their evaluations on the new pool of 12 candidate generative models.
+We also report attribution accuracy of the MFD methods, obtained by reproducing their evaluations on the pool of 12 candidate generative models.
 The remaining columns indicate whether the method is differentiable, whether it admits an analytic approximation, and which attack strategies are applicable.
@@
 We train three neural network models to evaluate fingerprint robustness across different attack settings.
-All models are optimized with Adam for 100 epochs on $1,000$ images/fingerprints per generative model, using a batch size of 32, a validation split of 20%, and early stopping to prevent overfitting.
+All models are optimized with Adam for 100 epochs on 1,0001{,}000 images/fingerprints per generative model, using a batch size of 32, a validation split of 20%, and early stopping to prevent overfitting.
 The learning rate is fixed at $1\times 10^{-4}$.
@@
 This model maps fingerprint features to source model labels and serves as the primary target for attacks.
-It is implemented as a MLP with hidden dimensions [512, 256, 128] and ReLU activations, trained with cross-entropy loss.
+It is implemented as an MLP with [512, 256, 128] hidden dimensions and ReLU activations, trained with cross-entropy loss.
 The architecture is fixed across all MFD methods, while the input dimensionality depends on the fingerprint size of the specific method.
@@
 We implement five adversarial attack strategies across the white-box (W1–W3) and black-box (B1–B2) threat models detailed in Section IV to evaluate fingerprint robustness.
-While removal attack are not targeted, forgery attacks require the attacker to induce a misclassification towards a specific target candidate generator label different from the true label.
+While removal attacks are not targeted, forgery attacks require the attacker to induce a misclassification towards a specific target candidate generator label different from the true label.
 In the evaluation, we choose the target label uniformly at random on each execution of a forgery attack.
@@
 The perturbation is bounded by $\epsilon=0.025$ (i.e., $\ell_{\infty}$ norm).
-We adopt the $\ell_{\infty}$ norm for its computational efficiency and its direct interpretation as a pixel value bound, while also verifying that perturbations remain visually acceptable.
+We adopt the $\ell_{\infty}$ norm for its computational efficiency and its direct interpretation as a pixel value bound, while allowing to verify that perturbations remain visually acceptable.
 Across all attacks, this $\epsilon$ level consistently yields LPIPS <0.05<0.05 and PSNR >35>35 dB, which indicates satisfactory perceptual quality [zhang2018lpips, huynh2008psnr] (see Appendix A for more details).
@@
 All gradient-based attacks support both removal and forgery objectives.
-For non-gradient attacks (B2), we employ four standard image perturbations with parameter settings chosen to preserve high image quality comparable to gradient-based attacks (i.e.
-LPIPS <0.05<0.05 and PSNR >35>35 dB).
+For non-gradient attacks (B2), we employ four standard image perturbations with parameter settings chosen to preserve high image quality comparable to gradient-based attacks (i.e., LPIPS <0.05<0.05 and PSNR >35>35 dB).
 Specifically, we apply additive Gaussian noise with $\sigma=0.005$ (B2-Noise), Gaussian blur with $\sigma=0.5$ using a $3{\times}3$ kernel (B2-Blur), JPEG compression at quality $95$ (B2-JPEG), and isotropic downsampling to $0.9\times$ followed by upsampling (B2-Resize).
-Each perturbation is applied once independently to the original image, simulating one single adversarial query.
+Each perturbation is applied once independently to the original image, each simulating one single removal attempt.
 We also conduct an ablation over these hyperparameters and report their impact on attack success in Figure 9 in Appendix D.
@@
 The reported results are the mean and standard deviation over $5$ independent runs, where each run involves training all models in the pipeline and then conducting an attack.
-Upon acceptance, we will publish all our source code with documentation on how to reproduce the experiments.
 VI Results
 This section presents the results of our security evaluation.
-We organize it into five parts: first, we assess MFD technique robustness to removal across threat models and attack strategies; second, we repeat the evaluation for forgery; third, we study the relationship between removal and forgery; fourth, we test whether higher utility of the MFD technique is an indicator of vulnerability to the attacks; and, fifth, we investigate the impact of model’s type and architecture on ASR.
+We organize it into five parts: first, we assess MFD technique robustness to removal across threat models and attack strategies; second, we repeat the same evaluation for forgery; third, we study the relationship between removal and forgery; fourth, we test whether higher utility of the MFD technique is an indicator of attack success; and, fifth, we investigate the impact of the model’s type and architecture on ASR.
 Overall, our evaluation reveals that existing MFD techniques are highly susceptible to adversarial manipulation.
@@
 The maximum values per method are highlighted in bold.
-The row separators indicate the domains in Table I: RGB, Frequency, and Learned.
+The row separators indicate the domains: RGB, Frequency, and Learned.
 Method W1 W2 W3 B1 B2-Noise B2-Blur B2-JPEG B2-Resize
-McCloskey18 [McCloskey18] 36.70 $\pm$ 5.16 59.36 $\pm$ 28.81 13.49 $\pm$ 0.55 2.09 $\pm$ 1.00 13.33 $\pm$ 5.70 2.88 $\pm$ 1.75 11.97 $\pm$ 5.89
-Nataraj19 [Nataraj19] 97.53 $\pm$ 1.84 86.54 $\pm$ 12.49 84.75 $\pm$ 0.94 27.97 $\pm$ 0.85 38.31 $\pm$ 3.16 46.16 $\pm$ 4.09 51.42 $\pm$ 1.14
-Nowroozi22 [Nowroozi22] 84.94 $\pm$ 1.32 67.87 $\pm$ 1.75 84.55 $\pm$ 0.97 22.47 $\pm$ 1.69 29.00 $\pm$ 1.26 32.43 $\pm$ 1.66 42.39 $\pm$ 2.87
-Song24-RGB [Song24] 91.38 $\pm$ 2.63 2.34 $\pm$ 0.61 0.00 $\pm$ 0.00 1.19 $\pm$ 0.40 0.18 $\pm$ 0.32 0.00 $\pm$ 0.00
-Marra19a [Marra19a] 100.00 $\pm$ 0.00 40.71 $\pm$ 5.20 9.55 $\pm$ 2.10 1.67 $\pm$ 1.07 13.50 $\pm$ 9.26 4.10 $\pm$ 1.06 5.04 $\pm$ 1.18
-Durall20 [Durall20] 99.90 $\pm$ 0.20 95.23 $\pm$ 7.93 99.46 $\pm$ 0.94 20.83 $\pm$ 5.80 61.63 $\pm$ 2.21 39.04 $\pm$ 4.51 94.12 $\pm$ 1.11
-Dzanic20 [Dzanic20] 98.90 $\pm$ 1.71 99.10 $\pm$ 1.37 99.78 $\pm$ 0.22 40.33 $\pm$ 0.18 90.59 $\pm$ 4.08 48.26 $\pm$ 0.54 98.63 $\pm$ 0.61
-Giudice21 [Giudice21] 100.00 $\pm$ 0.00 92.74 $\pm$ 8.01 21.99 $\pm$ 1.30 59.21 $\pm$ 3.61 45.48 $\pm$ 1.80 81.18 $\pm$ 1.76
-Corvi23-R [Corvi23] 100.00 $\pm$ 0.00 35.16 $\pm$ 3.84 2.44 $\pm$ 1.03 17.02 $\pm$ 4.26 1.11 $\pm$ 0.37 25.20 $\pm$ 6.53
-Corvi23-S [Corvi23] 77.44 $\pm$ 27.47 15.08 $\pm$ 7.42 0.42 $\pm$ 0.72 27.90 $\pm$ 18.07 0.42 $\pm$ 0.72 32.21 $\pm$ 15.68
-Song24-Freq [Song24] 31.08 $\pm$ 3.82 70.56 $\pm$ 6.94 5.30 $\pm$ 0.87 50.55 $\pm$ 3.78 6.79 $\pm$ 1.55 68.64 $\pm$ 0.63
-Qian20 [Qian20] 90.71 $\pm$ 6.43 82.05 $\pm$ 6.12 56.08 $\pm$ 38.06 67.12 $\pm$ 28.49 26.28 $\pm$ 6.65 75.90 $\pm$ 20.89
-Wang20 [Wang20] 100.00 $\pm$ 0.00 78.19 $\pm$ 18.90 0.67 $\pm$ 0.77 37.08 $\pm$ 54.53 34.17 $\pm$ 57.01 49.04 $\pm$ 44.23
-Song24-SL [Song24] 84.45 $\pm$ 0.85 55.86 $\pm$ 1.97 2.68 $\pm$ 0.74 20.18 $\pm$ 1.66 7.77 $\pm$ 1.51 34.61 $\pm$ 2.59
+McCloskey18 [McCloskey18] $36.70$ p m 5.16 $59.36$ p m 28.81 $13.49$ p m 0.55 $2.09$ p m 1.00 $13.33$ p m 5.70 $2.88$ p m 1.75 $11.97$ p m 5.89
+Nataraj19 [Nataraj19] $97.53$ p m 1.84 $86.54$ p m 12.49 $84.75$ p m 0.94 $27.97$ p m 0.85 $38.31$ p m 3.16 $46.16$ p m 4.09 $51.42$ p m 1.14
+Nowroozi22 [Nowroozi22] $84.94$ p m 1.32 $67.87$ p m 1.75 $84.55$ p m 0.97 $22.47$ p m 1.69 $29.00$ p m 1.26 $32.43$ p m 1.66 $42.39$ p m 2.87
+Song24-RGB [Song24] $91.38$ p m 2.63 $2.34$ p m 0.61 $0.00$ p m 0.00 $1.19$ p m 0.40 $0.18$ p m 0.32 $0.00$ p m 0.00
+Marra19a [Marra19a] $100.00$ p m 0.00 $40.71$ p m 5.20 $9.55$ p m 2.10 $1.67$ p m 1.07 $13.50$ p m 9.26 $4.10$ p m 1.06 $5.04$ p m 1.18
+Durall20 [Durall20] $99.90$ p m 0.20 $95.23$ p m 7.93 $99.46$ p m 0.94 $20.83$ p m 5.80 $61.63$ p m 2.21 $39.04$ p m 4.51 $94.12$ p m 1.11
+Dzanic20 [Dzanic20] $98.90$ p m 1.71 $99.10$ p m 1.37 $99.78$ p m 0.22 $40.33$ p m 0.18 $90.59$ p m 4.08 $48.26$ p m 0.54 $98.63$ p m 0.61
+Giudice21 [Giudice21] $100.00$ p m 0.00 $92.74$ p m 8.01 $21.99$ p m 1.30 $59.21$ p m 3.61 $45.48$ p m 1.80 $81.18$ p m 1.76
+Corvi23-R [Corvi23] $100.00$ p m 0.00 $35.16$ p m 3.84 $2.44$ p m 1.03 $17.02$ p m 4.26 $1.11$ p m 0.37 $25.20$ p m 6.53
+Corvi23-S [Corvi23] $77.44$ p m 27.47 $15.08$ p m 7.42 $0.42$ p m 0.72 $27.90$ p m 18.07 $0.42$ p m 0.72 $32.21$ p m 15.68
+Song24-Freq [Song24] $31.08$ p m 3.82 $70.56$ p m 6.94 $5.30$ p m 0.87 $50.55$ p m 3.78 $6.79$ p m 1.55 $68.64$ p m 0.63
+Qian20 [Qian20] $90.71$ p m 6.43 $82.05$ p m 6.12 $56.08$ p m 38.06 $67.12$ p m 28.49 $26.28$ p m 6.65 $75.90$ p m 20.89
+Wang20 [Wang20] $100.00$ p m 0.00 $78.19$ p m 18.90 $0.67$ p m 0.77 $37.08$ p m 54.53 $34.17$ p m 57.01 $49.04$ p m 44.23
+Song24-SL [Song24] $84.45$ p m 0.85 $55.86$ p m 1.97 $2.68$ p m 0.74 $20.18$ p m 1.66 $7.77$ p m 1.51 $34.61$ p m 2.59
 We evaluate all MFD methods against their compatible attack strategies, as summarized in Table I.
@@
 In many cases, the strongest attack achieves over $80\%$ ASR under the $\ell_{\infty}$ norm constraint, indicating that fingerprints can often be suppressed while maintaining high perceptual quality.
-Corresponding LPIPS and PSNR values remain within good quality ranges for most attacks, and representative attacked images confirm that the perturbations are imperceptible (see Tables IV, V, and Figure 5 in Appendix A).
-The main exceptions are Marra19a and Song24-RGB, which exhibit strong resistance to most attacks unless white-box access is available.
+Corresponding LPIPS and PSNR values remain within good image quality ranges for most attacks, and upon visual inspection of representative attacked images we confirm that the perturbations are imperceptible (see Tables IV, V, and Figure 5 in Appendix A).
+The exceptions are Marra19a and Song24-RGB, which exhibit strong resistance to most attacks unless white-box access is available.
 McCloskey18, Corvi23-R, and Corvi23-S also show some resistance to black-box attacks; however, their baseline (non-adversarial) accuracies are substantially lower.
@@
 White-box attacks (W1–W3) generally outperform black-box attacks (B1 and B2) across MFD techniques, which is expected since white-box attackers have privileged visibility into the MFD system.
-Among white-box, W1 attacks consistently achieve the best performance—when they are applicable.
+Among white-box attacks, W1 consistently achieves the best performance—when applicable.
 However, W1 attacks are not applicable for techniques that do not support direct gradient optimization.
@@
 In many applications, such as forensic use cases, MFD techniques cannot realistically be kept secret, making B1 removal attacks a practical threat.
-Although adaptive strategies (W1–W3, and B1) always outperform fingerprint-agnostic strategies based on simple image manipulations (B2), these generic black-box perturbations alone often achieve ASRs $50\%$ and achieve ASR $90\%$ for one of the techniques (Dzanic20), indicating that fingerprint traces can be significantly disrupted without any knowledge of the underlying MFD technique.
-This further demonstrates the fragility of current MFD methods and heightens concerns about the applicability of these techniques in adversarial settings where not even costly attacks are needed.
+Although adaptive strategies (W1–W3, and B1) always outperform fingerprint-agnostic strategies based on simple image manipulations (B2), these generic black-box perturbations alone often achieve ASRs $50\%$, with one of the techniques (Dzanic20) being particularly vulnerable (ASR $90\%$).
+This result indicates that fingerprint traces can be significantly disrupted without any knowledge of the underlying MFD technique, demonstrating the fragility of current MFD methods and raising concerns about the applicability of these techniques in adversarial scenarios.
 ASRs show low standard deviation across most attack instances.
@@
 difference in ASR.
-The MFD technique appears to be the primary driver of this variance, creating uncertainty for users about attack robustness.
+The choice of MFD technique appears to be the main driver of this variance, creating uncertainty for users about attack robustness.
 This variability suggests certain samples are inherently more vulnerable to fingerprint removal.
@@
 Method W1 W2 W3 B1
-McCloskey18 [McCloskey18] 7.51 $\pm$ 2.82 10.43 $\pm$ 2.81 2.10 $\pm$ 0.06
-Nataraj19 [Nataraj19] 51.64 $\pm$ 2.78 9.85 $\pm$ 3.17 11.27 $\pm$ 0.72
-Nowroozi22 [Nowroozi22] 12.09 $\pm$ 3.57 6.84 $\pm$ 0.73 10.20 $\pm$ 1.57
-Song24-RGB [Song24] 16.42 $\pm$ 4.59 0.00 $\pm$ 0.00
-Marra19a [Marra19a] 97.85 $\pm$ 2.90 4.46 $\pm$ 1.28 0.50 $\pm$ 0.52
-Durall20 [Durall20] 57.67 $\pm$ 8.63 23.49 $\pm$ 2.70 18.18 $\pm$ 1.13
-Dzanic20 [Dzanic20] 15.19 $\pm$ 5.09 25.53 $\pm$ 9.61 25.51 $\pm$ 3.93
-Giudice21 [Giudice21] 98.92 $\pm$ 0.39 13.95 $\pm$ 4.09
-Corvi23-R [Corvi23] 99.77 $\pm$ 0.40 3.67 $\pm$ 1.38
-Corvi23-S [Corvi23] 9.12 $\pm$ 4.06 2.52 $\pm$ 1.47
-Song24-Freq [Song24] 3.31 $\pm$ 1.38 6.66 $\pm$ 0.88
-Qian20 [Qian20] 18.12 $\pm$ 15.79 12.51 $\pm$ 1.64
-Wang20 [Wang20] 99.55 $\pm$ 0.77 5.68 $\pm$ 5.09
-Song24-SL [Song24] 10.09 $\pm$ 1.90 7.29 $\pm$ 1.47
+McCloskey18 [McCloskey18] $7.51$ p m 2.82 $10.43$ p m 2.81 $2.10$ p m 0.06
+Nataraj19 [Nataraj19] $51.64$ p m 2.78 $9.85$ p m 3.17 $11.27$ p m 0.72
+Nowroozi22 [Nowroozi22] $12.09$ p m 3.57 $6.84$ p m 0.73 $10.20$ p m 1.57
+Song24-RGB [Song24] $16.42$ p m 4.59 $0.00$ p m 0.00
+Marra19a [Marra19a] $97.85$ p m 2.90 $4.46$ p m 1.28 $0.50$ p m 0.52
+Durall20 [Durall20] $57.67$ p m 8.63 $23.49$ p m 2.70 $18.18$ p m 1.13
+Dzanic20 [Dzanic20] $15.19$ p m 5.09 $25.53$ p m 9.61 $25.51$ p m 3.93
+Giudice21 [Giudice21] $98.92$ p m 0.39 $13.95$ p m 4.09
+Corvi23-R [Corvi23] $99.77$ p m 0.40 $3.67$ p m 1.38
+Corvi23-S [Corvi23] $9.12$ p m 4.06 $2.52$ p m 1.47
+Song24-Freq [Song24] $3.31$ p m 1.38 $6.66$ p m 0.88
+Qian20 [Qian20] $18.12$ p m 15.79 $12.51$ p m 1.64
+Wang20 [Wang20] $99.55$ p m 0.77 $5.68$ p m 5.09
+Song24-SL [Song24] $10.09$ p m 1.90 $7.29$ p m 1.47
 VI-B Robustness to Forgery
 Table III shows that forgery ASRs are below those of removal across the board.
-Forgery is a fundamentally harder task than removal: with 12 models in the attribution pool, successful forgery requires targeted misattribution to a specific model, while removal only requires misclassification into any incorrect class.
-Geometrically, removal corresponds to crossing the nearest decision boundary, whereas forgery demands reaching a potentially narrow and distant target region.
-When only two models are considered (which is the case for general deepfake detection tasks), the two objectives coincide and removal corresponds to forgery, but as the number of models grows, the decision landscape becomes increasingly complex and harder to manipulate by a forgery attack.
+Forgery is a fundamentally harder task than removal: with 12 models in the pool, successful forgery requires targeted misattribution to a specific model, while removal succeeds upon misclassification into any incorrect class.
+Geometrically, removal corresponds to crossing the nearest decision boundary, whereas forgery requires pushing the image to a potentially farther target region.
+When MFD considers only two classes, such as in deepfake detection, removal and forgery become equivalent: erasing traces of one source implies attribution to the other.
+But as the number of classes grows, the decision landscape becomes increasingly complex, making forgery attacks harder to execute.
 Despite this difficulty, many methods still exhibit substantial forgery ASRs, especially in white-box settings.
@@
 In practice, the attacker may be interested in a specific model, or a specific model family that might be more vulnerable to attacks, in which case the average ASR within the family might be significantly higher.
-Furthermore, since our gradient-based attacks are generic and not tailored to the internal design of each method, method-specific adaptations could plausibly yield even higher forgery success rates.
+Furthermore, our gradient-based attacks are method-agnostic rather than exploiting the specific architecture of each detector; tailored attacks would likely achieve even higher forgery ASRs.
 Similar to removal attacks, forged samples cannot be visually distinguished from their original counterparts, maintaining an LPIPS <0.05<0.05 and PSNR >35>35 dB across all forged samples (see Tables VI, VII in Appendix A).
-It is important to note that higher ASR for forgery and removal could be achieved if the adversary relaxed our strict $\ell_{\infty}$ bound constraints, as explained in Appendix D.
+It is important to note that even higher ASRs for both forgery and removal are feasible if the adversary relaxes the constraint on image quality ($\ell_{\infty}$ bound), as explained in Appendix D.
 Recall that we discarded B2 attack strategies from the forgery attack evaluation because they lack specificity to embed a targeted model fingerprint into the forged sample.
@@
 In contrast, ASR variance for McCloskey18 is relatively lower than in removal attacks, while Qian20 now exhibits the highest variance.
-This indicates that while the objectives of removal and forgery are similar, they are not perfectly complementary—a finding that motivates a deeper investigation of their relationship.
+This finding indicates that while the objectives of removal and forgery are similar, they are not perfectly complementary, and motivates a deeper investigation of the relationship between removal and forgery.
 VI-C Removal vs. Forgery
@@
 2, MFD methods that are more susceptible to removal are also generally more vulnerable to forgery.
-This trend is intuitive: since all reported attacks are gradient-based, if gradients can efficiently guide perturbations toward a distant forgery region, they can likewise push samples outside their current decision region, enabling removal.
+This trend is intuitive: since all reported attacks are gradient-based, if gradients can efficiently guide perturbations toward a distant forgery region, they can likewise push samples outside their current decision region, leading to removal.
 Figure 2: Correlation between removal and forgery attack vulnerabilities across MFD methods.
@@
 Green circles indicate outliers with better trade-off.
-See Figure 7 in Appendix C for individual whit-box attack data.
+See Figure 7 in Appendix C for a more detailed analysis on white-box attacks.
 However, for the black-box scenario, this correlation does not imply that high removal ASR translates into high forgery ASR in absolute value terms.
@@
 In this setting, removing fingerprints of one model effectively forces the sample to be attributed to the other, making removal and forgery effectively the same goal.
-By contrast, as the number of models in the pool grows, removal remains comparably feasible—since pushing samples outside the decision region of the true model does not fundamentally depend on the pool size and other models—while forgery can become increasingly difficult.
-This is because forging requires moving a sample into a specific target region, and the difficulty of reaching such regions grows with the number and arrangement of models’ decision boundaries.
-Hence, both removal and forgery ASR scales can be affected by the geometric landscape of decision regions across models in each individual model attribution task.
+As the model pool grows, removal remains comparably feasible because pushing samples outside the true model’s decision region does not depend on the other models.
+Forgery, by contrast, becomes increasingly difficult: it requires moving a sample into a specific target region, and reaching that region becomes harder with more decision boundaries.
+Both removal and forgery ASR scales can be affected by the geometric landscape of decision regions across models in each individual model attribution task.
 The precise relationship between removal and forgery thus depends on the size of the model pool and the characteristics of the models.
@@
 As a result, they offer limited utility for attribution tasks, despite the strong detection performance originally reported in the deepfake setting.
-To examine the trade-off between attribution accuracy and adversarial robustness for MFD methods, we plot this relationship for both removal and forgery attacks in Figure 3 under white-box and black-box scenarios respectively.
-Broadly, we observe that the utility–robustness tension is pronounced for both removal and forgery attacks: methods with higher attribution accuracy tend to suffer higher ASR, indicating weaker robustness.
+To examine the trade-off between attribution accuracy and adversarial robustness for MFD methods, we plot in Figure 3 this relationship for both removal and forgery attacks under white- and black-box scenarios.
+Broadly, we observe a utility–robustness tension for both removal and forgery attacks: methods with higher attribution accuracy tend to suffer higher ASR, indicating weaker robustness.
 Most high-utility methods (accuracy >0.7>0.7) exhibit significantly higher vulnerability than low-utility ones.
-For example, neural network–based methods such as Wang20 achieve the highest attribution accuracy; yet, it is also among the most fragile methods to both removal and forgery, particularly under white-box attacks.
+For example, neural network–based methods such as Wang20 achieve the highest attribution accuracy but are also among the most vulnerable to both removal and forgery, particularly under white-box attacks.
 Nonetheless, there are notable exceptions: for instance, Marra19a and Song24-RGB under black-box attacks, and Song24-Freq under white-box attacks.
-These methods pair moderate attribution accuracy with substantially stronger robustness, making them stand out as more balanced candidates among the evaluated approaches.
+These methods pair moderate attribution accuracy with substantially stronger robustness, making them stand out as the most balanced candidates among all the evaluated approaches.
 Overall, these findings highlight a fundamental trade-off: methods with high attribution accuracy often lack robustness, whereas more robust methods typically sacrifice accuracy.
-Promising directions for future work include: (i) improving residual- and manifold-based approaches such as Marra19a and Song24 to boost accuracy; (ii) strengthening CNN-based methods like Wang20 against adversarial perturbations; and (iii) exploring hybrid or ensemble designs that combine the robustness of residual/manifold methods with the strong attribution accuracy of CNN-based approaches.
+The results indicate a few directions for future work, including: (i) improving residual- and manifold-based approaches such as Marra19a and Song24 to boost accuracy; (ii) strengthening CNN-based methods like Wang20 against adversarial perturbations; and (iii) exploring hybrid or ensemble designs that combine the robustness of residual/manifold methods with the strong attribution accuracy of CNN-based approaches.
 VI-E Attack Variance Across Models
 Figure 4: Model vulnerability per MFD method, averaged over all applicable attack types.
-Each subplot plots removal ASR (x) vs.
-forgery ASR (y) for the generative models (indicated by colors and markers) tied to that method; ASR spans 0–1 and error bars show standard deviation across different attack types.
-The final subplot averages the data over all MFD methods (with standard deviation representing statistics over different MFD methods), showing general vulnerability trend of generative models - models with more resilience to removal are more vulnerable to forgery, and vice versa.
+Each subplot shows removal ASR ($x$) vs.
+forgery ASR ($y$) for the generative models (indicated by colors and markers) tied to that method; error bars show standard deviation across different attack types.
+The final subplot averages the data over all MFD methods (with standard deviation representing statistics over different MFD methods), showing a general vulnerability trend of generative models: models with more resilience to removal are more vulnerable to forgery, and vice versa.
 To examine whether particular generative models are inherently more vulnerable to certain adversarial attacks, we evaluated per-model vulnerability across all MFD methods.
-Figure 4 plots forgery ASR versus removal ASR for each model under each method, averaging over all applicable attack strategies.
-We define a model’s removal vulnerability as how easily its own images can be perturbed to be misclassified as coming from any other model, and its forgery vulnerability as how easily images from other models can be perturbed to be misclassified as coming from this model.
-The results show that vulnerability is not consistently high or low for any given model or model family.
+Figure 4 plots forgery ASR versus removal ASR for each model under each MFD method, averaging over all applicable attack strategies.
+The results show that attack vulnerability is not consistently high or low for any given model or model family.
 Instead, it varies strongly with the MFD method.
@@
 In forgery, ADM is significantly more vulnerable under Nataraj19 and Nowroozi22, yet this pattern does not generalize across all MFD methods.
-Despite this variability, an important and obvious trend emerges: across MFD methods, image generation models that are more robust to removal tend to be more vulnerable to forgery, and vice versa.
+Despite this variability, a trend emerges: across MFD methods, image generation models that are more robust to removal tend to be more vulnerable to forgery, and vice versa.
 This trade-off is visible in the final subplot of Figure 4, and is also particularly clear under methods such as Nowroozi22, Song24-Freq, and Qian20.
@@
 VII Discussion
-Researchers have developed MFD techniques to provide provenance of model outputs and thus address some of the accountability challenges posed by today’s increasingly powerful generative models.
-In this paper, we provide the first systematic security analysis of these techniques, which includes over 14 MFD techniques across a range of possible threat models and attack strategies.
-Despite the reported high accuracy of these techniques in a clean setting in the literature, our findings reveal that they are highly vulnerable to adversarial manipulations.
-Our results show that most techniques fail to provide reliable model attribution under removal attacks.
-For most MFD techniques, our attacks remove model fingerprints across various access assumptions without visually degrading the attacked images.
-Most concerning, simple black-box image perturbations succeed in removing fingerprints, with attacks achieving over $90\%$ success for some techniques.
-These findings indicate that fingerprint traces can be disrupted without sophisticated attacks or knowledge of the underlying MFD system.
-Implications for MFD Technique Deployment.
-These results fundamentally undermine the applications proposed in the academic literature.
-The stark contrast between MFD’s high accuracy in the clean setting (often $90\%$) and its fragility under attack, uncovers a gap that can lead to a false sense of security if unacknowledged.
-For example, such high attack success rates imply that investigators and model providers cannot use these techniques to reliably trace images to the source models, compromising their forensic value and undermining the accountability they are intended to provide.
-Importantly, while forgery attacks are significantly less successful than removal, effective removal alone severely compromises MFD’s value for such applications.
-Even in the clean setting, existing evaluations may have overestimated MFD performance for model attribution.
-Current studies evaluate small closed worlds of at most 20 different models, raising questions about how representative they are of real-world scenarios.
-Practical deployments could be far more complex.
-For example, detecting the source of an arbitrary image would require distinguishing among orders of magnitude more models.
-Furthermore, real-world scenarios might involve highly similar models: the ML community often reuses public datasets for training and relies on open sourced learning algorithms, potentially producing similar model fingerprints—a situation not represented by the small, diverse model pools considered in the literature.
-Conversely, other practical scenarios might involve smaller model pools.
-Additional sources of evidence can narrow down the candidate pool to one closer to those in the literature, potentially enabling effective MFD deployment.
-Therefore, we encourages future papers presenting new MFD schemes to evaluate them in realistic open worlds.
-However, the success of our attacks implies that these techniques remain unreliable even in such ideal conditions, highlighting the urgent need for developing MFD techniques that are robust in adversarial environments.
-Implications for the Adversarial Arms-Race.
-Our results also have implications for the strategies chosen by attackers and fingerprinters.
-Since attackers cannot predict which MFD technique will be applied, they must develop attacks capable of removing fingerprints across all MFD techniques.
-This situation might seem like a whack-a-mole game that favors the fingerprinter, who can apply multiple techniques simultaneously.
-However, our results add nuances to this: we observe a tension between utility and robustness, where the most robust methods (e.g., Marra19a, Song24-RGB) exhibit higher detection error than the more vulnerable ones.
-Consequently, even when the fingerprinter deploys all 14 techniques, the attacks increase overall attribution uncertainty and thus undermine value of these techniques for forensic applications.
-Limitations and Future Work.
-One of the main limitations of our work is that we cannot conclude that the removal attacks completely erase the fingerprints: they supress enough signal to impair model attribution under a specific experimental setting, but there may be enough signal to still enable deepfake detection.
-Future work could further investigate the effectiveness of the attacks against MFD techniques for deepfake detection.
-Another avenue for future work is a more thorough exploration of the attack hyperparameter space.
-Although we tuned parameters that had significant impact, we made no attempt at optimization.
-Similarly, although future work could explore attacks tailored to specific techniques, attackers must be strategic: since attackers cannot predict which MFD techniques fingerprinters will deploy, attackers are better off increasing overall maximum success across techniques.
-Therefore, even highly successful bespoke attacks provide only marginal advantage if other techniques remain robust or if increased success against one method comes at the cost of reduced effectiveness against others.
-From the perspective of mitigations, an important direction for future work is to evaluate defenses from adversarial machine learning, such as adversarial training or randomized smoothing, applied to the attribution model.
-A thorough investigation would require adapting these defenses to handle both removal and forgery attacks and developing corresponding adaptive attacks to properly assess robustness.
-Such methods must also consider the typical utility–robustness trade-off, as improved robustness often comes at the cost of reduced clean accuracy.
-We view this as a substantial research effort that merits a dedicated study, and thus leave the systematic design and evaluation of defenses for future work.
-Additionally, we suspect that fingerprint suppression may introduce detectable patterns.
-We leave for future work investigating the detectability of the attacks.
-This capability would be particularly valuable for countering forgery attacks, as early detection could prevent adversaries from successfully misattributing content.
-Another avenue for defense is to explore hybrid approaches.
-In particular, despite their lower detection accuracies, our results identify Marra19a and Song24-RGB’s fingerprint features as particularly robust under black-box conditions.
-Future work could refine the features to maintain robustness while improving performance.
-Active vs.
-Passive Provenance Mechanisms.
-Watermarking has emerged as an alternative to fingerprinting for providing similar accountability guarantees in the deployment of generative models.
-The increased control of watermarking over the identifying signal makes active techniques more suitable for adversarial environments.
-Recently proposed watermarking schemes demonstrate robustness against sophisticated removal attempts while providing undetectability and formal security guarantees [yang2024gaussian, gunn2024undetectable].
-On the other hand, watermarking requires integration during model training or inference, which might introduce friction in its adoption by the industry.
-We encourage future work to further explore trade-offs between these two types of techniques or strengthen performance and robustness by combining the best of both worlds.
+This paper asks a practical question: how well does model fingerprint detection (MFD) hold up when the content creator actively tries to evade attribution?
+We address this by evaluating 14 MFD methods on a 12-model attribution task under three attacker knowledge levels (White-box, Black-box Access I, and Black-box Access II), two attacker goals (removal and forgery), and five attack strategies.
+Our results support three main conclusions.
+(1) Removal attacks are effective.
+For most MFD methods, the attacker can remove the fingerprints with imperceptible perturbations, even without information about the model or the MFD technique (Table II).
+(2) Forgery is relatively less effective but should not be dismissed.
+We still observe significant forgery success for many MFD methods, especially in white-box settings (Table III).
+(3) High clean accuracy does not imply robustness.
+Our results uncover a tension between clean attribution performance and adversarial robustness: none of the evaluated MFD methods is both consistently accurate and robust across the threat models we consider.
+VII-A Implications for Provenance Use-cases
+Many MFD papers motivate attribution as a provenance mechanism.
+Our findings suggest an important caveat: a method that performs well on clean benchmarks is unlikely to achieve the same performance in adversarial settings.
+This is especially relevant in forensic applications, where the subjects of an investigation have incentives to remove or forge fingerprints.
+Based on our evaluation, existing MFD methods are not yet ready for deployment in these contexts.
+A practical takeaway is that the evaluation methodology of MFD techniques in such applications should consider robustness, and should clearly define the threat model considered.
+It is also important to evaluate attribution under more realistic candidate sets.
+Real-world deployments would face large, open-world scenarios and frequent model updates that introduce closely related variants, such as fine-tuned checkpoints.
+Studying how attacks scale to realistic candidate pools is essential for understanding the practical effectiveness of MFD techniques, both with and without adversarial interference.
+VII-B What the Robustness Patterns Suggest
+The patterns in Tables II–III indicate that robustness is largely method-driven.
+Some MFD approaches appear relatively more resistant in constrained black-box settings, especially residual- and manifold-based MFD methods.
+However, this robustness is not universal and may not hold under white-box access.
+A useful way to view this is that black-box robustness can reflect limited attack transferability under a particular perturbation budget, rather than a fundamental inability to manipulate the fingerprint.
+This distinction matters operationally: if the attribution pipeline is public, it has leaked, or can be closely approximated, the threat of adaptive attacks becomes more relevant.
+These observations suggest a clear design objective for future research in MFD: methods should be developed and reported with an explicit robustness target, rather than focusing solely on clean accuracy.
+Based on the utility-robustness trade-off perspective, we identify two promising directions: (i) improving the clean accuracy of inherently robust representations, potentially leveraging NN-based classifiers to boost performance; and (ii) hardening high-accuracy fingerprints—such as MFD methods based on NN-learned features—against adaptive perturbations without sacrificing clean performance, potentially through techniques like adversarial training or randomized smoothing [cohen2019certified].
+VII-C Limitations and Future Work
+First, we examine multi-class attribution, a task inherently more complex than binary deepfake detection.
+A critical distinction lies in the relationship between attack types: while untargeted removal and targeted forgery are functionally equivalent in binary detection, they diverge significantly in the attribution setting.
+In multi-class attribution, untargeted removal is generally easier to achieve than evading detection of a deepfake because the adversary can exploit any of the multiple decision boundaries surrounding the correct class manifold.
+In contrast, targeted forgery remains more difficult, as it requires a precise shift into a specific, potentially distant region of the feature space.
+Consequently, an attack that successfully breaks attribution (removal) may still leave the sample within the broader manifold of synthetic content, failing to evade deepfake detection.
+Future work could study the gap between “detection-robustness” and “attribution-robustness” under the same threat model.
+Second, while our attacks are strong and cover multiple access levels, they do not span the full space of adaptive strategies.
+Future work could explore broader perturbation families and multi-step post-processing pipelines, and should also study whether fingerprint suppression or forgery leaves detectable traces that could be used as a warning signal that a verifier can flag for a more careful inspection.
+Third, the scope of our evaluation is constrained by the specific generator and dataset configuration (12 generators trained on FFHQ at $256\times 256$ resolution).
+Evaluating a broader range of generators and datasets, higher resolutions, and diverse types of distribution shifts would help quantify robustness under more diverse deployment conditions and thus provide a more complete understanding of MFD robustness.
+Finally, our results suggest that provenance systems should not rely on a single mechanism alone.
+Watermarking can provide a strong and unambiguous provenance signal when a watermark is embedded at generation time and remains detectable, and in such cases it is natural to rely primarily on the watermark.
+However, watermarking is often unavailable in practice: many images are generated by open-source or legacy models without watermarking, detection may require secret keys or vendor-side access, and real-world distribution or editing pipelines can remove or break watermark detectability altogether.
+Passive fingerprinting addresses a different need.
+It does not require cooperation at generation time and is detected post hoc, which allows provenance analysis even when watermarking is unavailable or not possible.
+Fingerprinting is therefore not a universal replacement for watermarking, nor is watermarking a replacement for fingerprinting.
+It is thus worthwhile exploring hybrid provenance systems that use watermarking as a high-confidence signal when present, and falls back to fingerprinting when it is not, while explicitly accounting for removal, forgery, and uncertainty under adversarial manipulation [yang2024gaussian, gunn2024undetectable].
 VIII Conclusion
-Model fingerprint detection (MFD) has been proposed as a means of providing provenance and accountability in the deployment of image generation models, with reported accuracies often exceeding $90\%$ in non-adversarial evaluations.
-However, our findings show that such results offer an incomplete picture of their reliability in the adversarial settings where attribution mechanisms are most critical.
-Through a systematic evaluation of 14 MFD techniques against 5 attack strategies, we expose a critical security gap: adversaries can reliably suppress or forge fingerprints, undermining the practical utility of current MFD methods for real-world attribution.
-While industry adoption of attribution mechanisms is already underway, it is unlikely to pause until robust solutions emerge.
-This urgency underscores the need not only to advance more resilient MFD techniques but also to develop complementary or alternative accountability mechanisms that remain effective under adversarial conditions.
-Looking forward, we identify several promising directions for strengthening model provenance of AI-generated images, including but not limited to leveraging more inherently robust methods such as residual- or manifold-based approaches, adversarially robust training of MFD systems, and hybrid designs that combine complementary strengths across existing model attribution techniques such as model watermarking.
+Model fingerprint detection has been proposed as a scalable means of providing provenance and accountability for AI-generated images, with many methods reporting strong attribution accuracy in non-adversarial evaluations.
+In this paper, we provide the first systematic security evaluation of such techniques for model attribution, formalizing three adversary knowledge levels and benchmarking 14 representative methods against five attack strategies under two attacker objectives: fingerprint removal and fingerprint forgery.
+Our results reveal a critical gap between clean and adversarial performance.
+Fingerprint removal is highly effective for most methods through small, imperceptible perturbations, even in black-box regimes.
+Forgery, in contrast, is less effective but can still succeed in black-box settings and achieves high success under white-box access on some methods.
+We also identify a utility–robustness trade-off: techniques with the highest clean attribution accuracy are frequently among the most vulnerable, and no evaluated method achieves both high accuracy and high robustness across all threat models.
+These findings suggest that current passive fingerprinting approaches alone are insufficient for high-stakes provenance in adversarial settings.
+Advancing reliable attribution will likely require (i) more robust fingerprint representations (e.g., residual/manifold-based), (ii) incorporating security considerations in the development and evaluation of new techniques, and (iii) hybrid designs that combine complementary provenance mechanisms, including active approaches like watermarking.
 Usage of LLMs
@@
 Acknowledgments
+We thank our anonymous reviewers for their valuable feedback, which substantially improved the paper.
 This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.
 Access to EIDF was facilitated through the University of Edinburgh’s Generative AI Laboratory GAIL Fellow scheme.
+Marc Juarez is a GAIL fellow and the recipient of a Google Research Scholar Award in Security.
 Appendix A Image Quality Under Attacks
@@
 As an example, Figure 5 provides a comprehensive visualization of successful removal attacks against the Song24-Freq MFD method across different generative models and attack strategies.
-The grid structure clearly illustrates how various attack types (W3, B1, and B2 with different perturbation techniques) perform on different target models, with each cell showing the original image, the perturbed result, and the scaled difference map.
+The grid structure clearly illustrates how various attack types (W3, B1, and B2 with different perturbation techniques) perform on different target models, with each cell showing the original image, the perturbed result, and the difference, which has been scaled to improve visibility.
 This visualization demonstrates the effectiveness of different attack strategies in removing model fingerprints while preserving the visual quality of the generated images.
-The difference images, scaled for visibility, reveal the spatial distribution of perturbations introduced by each attack method.
+The difference images show the spatial distribution of perturbations introduced by each attack.
 Figure 5: Visualization of successful removal attacks against the Song24-Freq MFD method for selected models.
@@
 The visualization demonstrates the effectiveness of various attack strategies in removing model fingerprints while maintaining visual quality.
-An “N/A” entry indicates non-existence of a successfully attacked image for the case.
+An “N/A” value indicates that there is no successfully attacked image for that entry.
 Appendix B Removal ASR vs Forgery ASR
-Here in Figure 6 we report the relationship between removal-attack ASR and forgery-attack ASR for each gradient-based attack type separately, rather than aggregating them using the best white-box ASR as done in the main text (i.e.
-Figure 2).
+In Figure 6 we report the relationship between removal ASR and forgery ASR for each gradient-based attack type separately, rather than aggregating them using the best white-box ASR as done in the main text (Figure 2).
 Specifically, we present results for W1, W2, and W3 individually.
 Note that certain attack types are only applicable to specific fingerprinting methods (e.g., W1 requires differentiability and is therefore only compatible with MFD methods such as Giudice21, Corvi23-R, Corvi23-S, Qian20, and Wang20).
-Importantly, the strong positive correlation between removal and forgery ASR persists for each attack type considered independently, reinforcing the robustness of the trend reported in the main paper.
+Importantly, the strong positive correlation between removal and forgery ASR persists for each independent attack type, reinforcing the trend reported in the main text.
 Figure 6: Correlation between removal and forgery attack vulnerabilities across MFD methods.
@@
 Appendix C Utility-Robustness Trade-off
-Here in Figure 7 we report the trade-off relationship between utility and attack ASR for each gradient-based attack type separately, rather than aggregating them using the best white-box ASR as done in the main text (i.e.
-Figure 3).
+In Figure 7 we report the trade-off between utility and attack ASR for each gradient-based attack type separately, rather than aggregating them using the best white-box ASR as done in the main text (Figure 3).
 Specifically, we present results for W1, W2, and W3 individually.
 Note that certain attack types are only applicable to specific fingerprinting methods.
-Figure 7: Trade-off between attribution utility (accuracy of the classifier $h$, y-axis) and robustness (ASR, x-axis).
-We plot the data with regard to each attack type separately in this figure, as opposed to Figure 3.
+Figure 7: Trade-off between attribution utility (accuracy of classifier $h$, $y$-axis) and robustness (ASR, $x$-axis).
+We plot the data with regard to each attack type separately, as opposed to Figure 3.
 Each point is an MFD method; error bars show standard deviation across 5 independent runs.
-Methods nearer the top-left deliver stronger utility–robustness trade-offs.
+Methods nearer the top-left achieve a better compromise between utility and robustness.
 Appendix D Hyperparameter Tuning
 We conducted hyperparameter tuning for several attack cases with respect to their strength parameters to illustrate the trade-off between ASR and image quality degradation.
-Specifically, we report ablation results for multiple attack cases using MFD Dzanic20 as a representative example, as shown in Figures 8 and 9.
+Specifically, we report ablation results for multiple attack cases using Dzanic20 as a representative example, as shown in Figures 8 and 9.
 For PGD-based attacks (W3 and B1), we varied the perturbation budget $\epsilon$ (the maximum allowed change per pixel) over the range {0.001, 0.005, 0.01, 0.05, 0.1, 0.5, 1.0}.
@@
   • Resize scaling factor: {0.9, 0.8, 0.7, 0.6, 0.5}.
-Figures 8 and 9 plot ASR and image quality indicators (LPIPS/PSNR) as functions of these attack-strength hyperparameters.
+Figures 8 and 9 plot ASR and image quality indicators (LPIPS/PSNR) as functions of these hyperparameters.
 As expected, increasing attack strength generally yields higher ASR, but at the cost of noticeably degraded image quality, particularly for PGD-based attacks.
+Figure 8: Effect of perturbation strength (PGD’s $\epsilon$) on attack success and image quality for removal (W3, B1) and forgery (B1) attacks.
+The curves show that there is a trade-off between robustness and image quality regulated by perturbation strength.
 […]
-[…]